### v2

# references.bib

Abstract
We propose the Block Coordinate Descent Network Simplex (BCDNS) method for solving large-scale discrete Optimal Transport (OT) problems. BCDNS integrates the Network Simplex (NS) algorithm with a block coordinate descent (BCD) strategy, decomposing the full problem into smaller subproblems per iteration and reusing basis variables to ensure feasibility. We prove that BCDNS terminates in a finite number of iterations with an exact optimal solution, and we characterize its per-iteration complexity as $O(sN)$, where $s\in(0,1)$ is a user-defined parameter and $N$ is the total number of variables. Numerical experiments demonstrate that BCDNS matches the classical NS method in solution accuracy, reduces memory footprint compared to Sinkhorn algorithm, achieves speed-ups of up to tens of times over the classical NS method, and exhibits runtime comparable to high-precision Sinkhorn algorithm.
Keywords: Optimal Transport, Network Simplex, Block Coordinate Descent, Sinkhorn, Wasserstein Distance
1.   Introduction
The Optimal Transport (OT) problem is to find the most efficient way to transport resources between a finite number of locations while minimizing transportation costs. It is rooted in mathematical optimization and has a rich history dating back to the 18th century. Gaspard Monge [monge1781deblais] first proposed it in 1781. Then in the mid-20th century, Leonid Kantorovich [kantorovich1942translocation] extended Monge’s framework to encompass continuous distributions, reformulating it as a linear programming model and significantly broadening its theoretical foundation and practical applications. In this thesis, we focus specifically on solving the discrete OT problem.
The OT problem has found applications in various disciplines. In machine learning, the OT problem has been used for tasks such as domain adaptation [courty2016optimal] and generative modeling  [arjovsky2017wasserstein]. In economics, the OT problem has provided insights into resource allocation, market equilibrium, and pricing mechanisms [galichon2016optimal]. In biology, the OT problem has been applied to analyze genetic expression patterns [schiebinger2019optimal].
The OT problem has found applications in various disciplines. In machine learning, the OT problem has been used for tasks such as domain adaptation [courty2016optimal] and generative modeling [arjovsky2017wasserstein]. In economics, the OT problem has provided insights into resource allocation, market equilibrium, and pricing mechanisms [galichon2016optimal]. In biology, the OT problem has been applied to analyze genetic expression patterns [schiebinger2019optimal]. In geophysics and signal processing, the OT problem has been employed for seismic waveform comparison and inversion. [liao2022fast].
Several methods have been developed to solve the OT problem. Each method has distinct strengths and limitations. The Hungarian method is a classical approach for solving linear assignment problems and can also be applied to OT problems [kuhn1955hungarian]. It achieves exact optimal solutions by constructing a cost matrix and iteratively reducing reducing dimensions of the matrix. However, its time complexity is $O(N^{2})$, where $N$ is the number of variables in OT problem, and hence it is not applicable for the large-scale OT problem. The Sinkhorn algorithm is a popular method for the large-scale OT problem due to its computational efficiency [cuturi2013sinkhorn]. It solves the OT problems with regularization of entropy, which is an approximation of the OT problem. Therefore it provides approximate solutions of the OT problem, and it may not be suitable for applications that require exact solutions. In recent years, many improvements based on the Sinkhorn algorithm have been proposed, such as the Greenkhorn algorithm [altschuler2017Near] and the Adaptive Primal–Dual Accelerated Gradient Descent (APDAGD) algorithm [dvurechensky2018Computational]. However, these methods still operate within the Sinkhorn framework and do not overcome the limitation of producing only approximate OT solutions. In contrast, the Network Simplex (NS) method, a specialized variant of the simplex method, is well known to yield exact solutions. Michael D. Grigoriadis [grigoriadis1986efficient] reported that the NS method is reliable among several algorithms. However, the NS method is not a polynomial time algorithm and it takes $O(N)$ time in each iteration, which leads to slow computational performance when applied to large-scale problems. We note that the above three methods needs at least $O(N)$ space complexity. To overcome the difficulty of large space complexity for the OT problem, Yue Xie et al. [xie2023randomized] proposed the Block Coordinate Descent (BCD) method for solving the OT problem. Their method constructs a small subproblem by randomly selecting variables based on predefined rules and solves it in each iteration. Since the subproblem is small, the BCD method does not require the O(N) space complexity. However, since the BCD method proposed by Yue Xie et al. adapts the random variable selection, it is not guaranteed to find an optimal solution within a finite number of iterations. Moreover, it sometimes takes a large number of iterations to get a reasonable approximate solution.
The motivation behind this research lies in the growing demand for efficient methods capable of solving large-scale OT problems with highly accurate solutions. As data sizes continue to grow, we recognize the increasing importance of solving large-scale OT problems with both speed and accuracy. To address these challenges, we aim to develop a method that enables fast computation of exact solutions for large-scale OT problems.
To achieve the goal of efficiently solving large-scale OT problems, we introduce the idea of the BCD method. The BCD method divides a large-scale OT problem into smaller subproblems, enabling a reduction in computational complexity. Although the BCD method is generally considered slow when exact solutions are required, we observe that the optimal solutions of the OT problem are typically sparse—due to the relatively small number of equality constraints compared to the number of variables—which significantly improves the efficiency of BCD. By employing the NS method to solve the subproblems, we guarantee that the final solution is also exact.
In this paper, we propose the Block Coordinate Descent Network Simplex (BCDNS) method, a novel algorithm that integrates the BCD and NS methods to achieve computational efficiency and solution accuracy. Our approach reduces computational complexity by decomposing large-scale OT problems into smaller subproblems, while the NS method ensures that each subproblem is solved with precision. Numerical experiments demonstrate significant performance gains of the BCDNS method over traditional approaches, addressing key limitations of existing methods.
In the BCD method, the selection of the block coordinate is crucial. Compared to existing BCD method [xie2023randomized] that randomly selects variables based on matrix properties, we propose that the BCDNS method first selects the basis variables from the previous subproblem as the initial solution for the next subproblem. This strategy ensures that the subproblem always has a feasible and an optimal solution. Moreover, by integrating the principles of the NS method, we ensure that the BCDNS method converges to the exact optimal solution within a finite number of computations. Furthermore, we categorize variables into two groups based on their likelihood of becoming optimal basis variables. This targeted selection process prioritizes variables with a higher probability of being optimal basis variables, thereby improving the efficiency of finding the optimal solution.
The outline of this paper is as follows. In Section 2, we introduce the foundational concepts of the OT problem and review several existing methods for solving it. In Section 3, we describe the core principles of the proposed BCDNS method. We first present its prototype along with the corresponding theoretical proofs. Then we develop a concrete implementation based on the prototype. Section 4 reports numerical experiments, including parameter estimation for the BCDNS method and a comparative performance analysis against existing approaches. Finally, Section 5 concludes the paper by summarizing the key contributions and discussing potential directions for future research.
2.   Discrete Optimal Transport
We consider a discrete OT problem [villani2009optimal] between $n$ supply locations and $m$ demand locations. Let $\{1,2,\dots,n\}$ denote the index set of supply points, and $\{1,2,\dots,m\}$ denote that of demand points.
Each supply point $i\in\{1,\dots,n\}$ provides a non-negative amount of mass $p_{i}\geq 0$, and each demand point $j\in\{1,\dots,m\}$ requires an amount $q_{j}\geq 0$. The total supply and demand are assumed to be balanced:
$\sum_{i=1}^{n}p_{i}=\sum_{j=1}^{m}q_{j}.$
Let $C=\{c_{ij}\}\in\mathbb{R}^{n\times m}$ be the cost matrix, where each entry $c_{ij}\geq 0$ denotes the cost of transporting one unit of mass from supply point $i$ to demand point $j$. A transport plan is represented by a matrix $X=\{x_{ij}\}\in\mathbb{R}^{n\times m}$, where $x_{ij}$ is the amount of mass transported from $i$ to $j$.
The discrete OT problem can then be formulated as the following linear program:
$\min_{X\in\mathbb{R}_{\geq 0}^{n\times m}}\quad$ $\sum_{i=1}^{n}\sum_{j=1}^{m}c_{ij}x_{ij}$ (1)
subject to $\sum_{j=1}^{m}x_{ij}=p_{i},\quad\forall i=1,\dots,n,$
$\sum_{i=1}^{n}x_{ij}=q_{j},\quad\forall j=1,\dots,m,$
$x_{ij}\geq 0,\quad\forall i=1,\dots,n,\;j=1,\dots,m.$
To express this in standard matrix-vector form, we define the following notation. Let $N=nm$ be the total number of transport variables, and let $M=n+m$ be the number of equality constraints. We vectorize the matrix $X$ into a column vector $\mathbf{x}\in\mathbb{R}^{N}$ by stacking its columns, i.e.,
$\mathbf{x}=\operatorname{vec}(X)=[x_{11},x_{21},\dots,x_{n1},x_{12},\dots,x_{nm}]^{\top}.$
We also define:
  • the cost vector $\mathbf{c}=\operatorname{vec}(C)\in\mathbb{R}^{N}$,
  • the supply vector $\mathbf{p}=[p_{1},\dots,p_{n}]^{\top}\in\mathbb{R}^{n}$,
  • the demand vector $\mathbf{q}=[q_{1},\dots,q_{m}]^{\top}\in\mathbb{R}^{m}$,
  • the constraint vector $\mathbf{b}=[\mathbf{p}^{\top},\,-\mathbf{q}^{\top}]^{\top}\in\mathbb{R}^{M}$.
  • The constraint matrix
    $A=\begin{bmatrix}I_{n}&I_{n}&\cdots&I_{n}\\ -1_{n}^{\top}&0&\cdots&0\\ 0&-1_{n}^{\top}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&-1_{n}^{\top}\end{bmatrix}\in\mathbb{R}^{M\times N},$
    where $I_{n}$ denotes the $n\times n$ identity matrix, and $1_{n}$ is an $n$-dimensional column vector of ones.
With this notation, we can rewrite the OT problem as:
$\min_{\mathbf{x}\in\mathbb{R}^{N}}\quad$ $\mathbf{c}^{\top}\mathbf{x}$ (2)
subject to $A\mathbf{x}=\mathbf{b},$
$\mathbf{x}\geq 0.$
3.   Existing Methods for Optimal Transport
In this section, we review several representative algorithms for solving the discrete OT problem. These include the classical NS method, the Sinkhorn algorithm based on entropy regularization which is widely used for large-scale problems, and the recently proposed BCD method.
3.1 Network Simplex Method for Optimal Transport
As shown in Appendix A, the constraint matrix $A\in\mathbb{R}^{M\times N}$ in the OT problem has rank $M-1$. This allows us to partition the set of variable indices $\mathcal{N}=\{1,\dots,N\}$ into two subsets: a basis index set $\mathcal{B}\subseteq\mathcal{N}$ with $|\mathcal{B}|=M-1$, and its complement $\mathcal{N}\setminus\mathcal{B}$. The columns of $A$ corresponding to $\mathcal{B}$, denoted $A_{B}\in\mathbb{R}^{M\times(M-1)}$, form a full-rank submatrix. The basic variables $\mathbf{x}_{B}$ can then be uniquely determined by the system $A_{B}\mathbf{x}_{B}=\mathbf{b}$, while the non-basic variables $\mathbf{x}_{N}$ are fixed at zero. A basic feasible solution further satisfies the non-negativity constraint $\mathbf{x}_{B}\geq 0$.
To determine whether a given basic solution is optimal, we examine its reduced-cost vector. Define the dual potential vector $\boldsymbol{\pi}\in\mathbb{R}^{M}$ via the dual feasibility condition $\boldsymbol{\pi}=(A_{B}^{\top})^{-1}\mathbf{c}_{B}$, where $\mathbf{c}_{B}$ is the cost subvector corresponding to the basis $B$. Then, for any index $k\in N\setminus B$, the reduced cost is computed as $r_{k}=c_{k}-\boldsymbol{\pi}^{\top}A_{k}$. Due to the structure of the OT problem, each column $A_{ij}$ has exactly one $+1$ and one $-1$, so the reduced cost simplifies to the closed-form expression
$r_{ij}=c_{ij}-(\pi_{i}-\pi_{n+j}).$
Optimality of the current basic solution is certified if all reduced costs are non-negative, i.e., $r_{ij}\geq 0$ for all $i,j$. If a variable $x_{i^{*}j^{*}}$ has $r_{i^{*}j^{*}}<0$, it indicates a direction of descent in the objective, and that variable is introduced into the basis.
To perform a pivot, we add the index $(i^{*},j^{*})$ to the basis $B$, inducing a unique cycle in the associated flow graph. Then determine the step size $\Delta$ by the minimum flow along the negative edges of the cycle, ensuring feasibility. Subsequently, update the flow: $x_{i^{*}j^{*}}\leftarrow\Delta$, and for each edge $(u,v)$ along the cycle, the value of $x_{uv}$ is incremented or decremented by $\Delta$ depending on the cycle direction. Finally, the variable that reaches zero and remove it from the basis, completing the pivot step [vanderbei2001linear].
Each iteration of the NS method involves computing the dual potential vector $\boldsymbol{\pi}$ in $O(M)$ time, evaluating all reduced costs in $O(N)$, and performing pivoting operations in $O(M)$. Since each pivot strictly improves the objective function by at least $-r_{i^{*}j^{*}}>0$, and the number of possible basic feasible solutions is finite, the NS method terminates in a finite number of iterations.
3.2 Block Coordinate Descent Method for Optimal Transport
In this subsection we introduce the BCD method for the OT problem in the LP form (2).
The BCD method iteratively selects a subset of coordinates $\mathcal{H}\subseteq\mathcal{N}=\{1,2,\dots,N\}$. For any feasible solution $\mathbf{x}$, we define a subproblem of the BCD method
$\𝒩D(\mathbf{x};\mathcal{H})\triangleq\arg\min_{\mathbf{d}\in\mathbb{R}^{N}}\left\{\mathbf{c}^{T}(\mathbf{x}+\mathbf{d})\mid\mathbf{x}+\mathbf{d}\geq 0,\;A\mathbf{d}=0,\;d_{i}=0,\;\forall i\in\mathcal{N}\backslash\mathcal{H}\right\}.$ (3)
$\𝒩q(\mathbf{x};\mathcal{H})\triangleq\min_{\mathbf{d}\in\mathbb{R}^{N}}\left\{\mathbf{c}^{T}(\mathbf{x}+\mathbf{d})\mid\mathbf{x}+\mathbf{d}\geq 0,\;A\mathbf{d}=0,\;d_{i}=0,\;\forall i\in\mathcal{N}\backslash\mathcal{H}\right\}.$ (4)
Here, $D(\mathbf{x};\mathcal{H})$ represents the set of optimal solutions for the subproblem with $\mathcal{H}$, and $q(\mathbf{x};\mathcal{H})$ denotes the optimal objective value corresponding to $d\in D(\mathbf{x};\mathcal{H})$ [tseng2009block].
The iterative BCD algorithm proceeds as follows:
  • Initialization: Start with a feasible solution $\mathbf{x_{0}}$ and an initial block $\mathcal{H}_{0}\subseteq\{1,\dots,N\}$.
  • Subproblem Optimization: For the current solution $\mathbf{x^{k}}$, solve the subproblem restricted to the working set $\mathcal{H}_{k}$ to find the descent direction $\mathbf{d^{k}}\in D(\mathbf{x^{k}};\mathcal{H}_{k})$. Update the solution as:
    $\mathbf{x^{k+1}}=\mathbf{x^{k}}+\mathbf{d^{k}}.$
  • Optimality Check: If optimality is achieved, terminate the algorithm. Else adjust the block $\mathcal{H}_{k}$ for the next iteration using predefined rules or adaptive heuristics.
The existing BCD method [xie2023randomized], randomly selects block coordinates based on certain properties of specific matrices. In [xie2023randomized], the authors propose the Accelerated Random Block Coordinate Descent algorithm and establish the following convergence theorem.
Theorem 1.
The gap between the current objective function value and the optimal objective function value converges with rate $1-v$, i.e.,
$\mathbf{c}^{\top}\mathbf{x}^{k+1}-\mathbf{c}^{\top}\mathbf{x}^{*}\leq(1-v)(\mathbf{c}^{\top}\mathbf{x}^{k}-\mathbf{c}^{\top}\mathbf{x}^{*}).$
Here, $v$ satisfies
$v\geq\zeta\frac{n(\tilde{p}-2)}{(n^{2}-3)(n!)^{2}},$
where $\zeta$ is a predefined probability parameter in the range $(0,1)$, and $\tilde{p}$ is a parameter representing the properties of the matrix in the range $[3,n]$.
These approaches provide a simple mechanism for constructing subproblems by selecting block coordinates. However, when the size n of the OT problems is large, $v$ becomes extremely small, which leads to slow convergence. Furthermore, these approaches do not account for the sparsity of the solution of the OT problem. Many variables that ultimately take zero values continue to be included in the block coordinates at each iteration. Moreover, these approaches have an inherent limitation: they cannot guarantee convergence to the optimal solution within a finite number of iterations. This limitation arises from the possibility of extreme cases where the randomly selected block of coordinates rare includes the optimal basis variables.
3.3 Sinkhorn Algorithm for Optimal Transport
Instead of solving the classical OT problem defined in (1), the Sinkhorn algorithm [cuturi2013sinkhorn] introduces an entropy regularization term
$\min_{X\in\mathbb{R}_{+}^{n\times m}}\;\sum_{i=1}^{n}\sum_{j=1}^{m}c_{ij}x_{ij}-\varepsilon H(X),$
where $\varepsilon>0$ is the regularization parameter, and $H(X)$ is the entropy function defined as
$H(X)=-\sum_{i=1}^{n}\sum_{j=1}^{m}x_{ij}\log x_{ij}.$
The parameter $\varepsilon$ controls the strength of the entropy regularization, balancing the trade-off between the classical transport cost and the entropy of the transport plan.
Introduce Lagrange multipliers $\alpha=(\alpha_{i})_{i=1}^{n}$ and $\beta=(\beta_{j})_{j=1}^{m}$ for the supply and demand constraints, respectively. The Lagrangian function of the regularized OT problem is
$\mathcal{L}(X,\alpha,\beta)=\sum_{i=1}^{n}\sum_{j=1}^{m}c_{ij}x_{ij}-\varepsilon H(X)+\sum_{i=1}^{n}\alpha_{i}\left(p_{i}-\sum_{j=1}^{m}x_{ij}\right)+\sum_{j=1}^{m}\beta_{j}\left(q_{j}-\sum_{i=1}^{n}x_{ij}\right).$
Taking the partial derivative of $\mathcal{L}$ with respect to $x_{ij}$ and setting it to zero, we obtain the optimality condition
$\frac{\partial\mathcal{L}}{\partial x_{ij}}=c_{ij}+\varepsilon(\log x_{ij}+1)-\alpha_{i}-\beta_{j}=0.$
Solving for $x_{ij}$, we find
$x_{ij}=\exp\left(\frac{\alpha_{i}+\beta_{j}-c_{ij}}{\varepsilon}-1\right).$
Let $u_{i}=e^{\alpha_{i}/\varepsilon}$ and $v_{j}=e^{\beta_{j}/\varepsilon}$, then the optimal transport plan can be expressed as
$x_{ij}=u_{i}e^{-c_{ij}/\varepsilon}v_{j}.$
Define $K=e^{-C/\varepsilon}$ as the matrix obtained by element-wise exponentiating the cost matrix $C=\{c_{ij}\}$. The transport plan $X$ can then be written as:
$X=\text{diag}(\mathbf{u})K\text{diag}(\mathbf{v}),$
where $\text{diag}(\mathbf{u})$ and $\text{diag}(\mathbf{v})$ are diagonal matrices whose diagonal entries are the elements of the vectors $\mathbf{u}$ and $\mathbf{v}$, respectively.
From the equality constraints $X\mathbf{1}_{m}=\mathbf{p}$ and $X^{\top}\mathbf{1}_{n}=\mathbf{q}$, we derive the following update rules
$\mathbf{u}^{(t+1)}=\frac{\mathbf{p}}{K\mathbf{v}^{(t)}},$
$\mathbf{v}^{(t+1)}=\frac{\mathbf{q}}{K^{\top}\mathbf{u}^{(t+1)}}.$
Here, the division is performed element-wise.
The Sinkhorn algorithm terminates when the change in the scaling factors $\mathbf{u}$ and $\mathbf{v}$ becomes sufficiently small
$\|\mathbf{u}^{(t)}-\mathbf{u}^{(t-1)}\|<\delta,\quad\|\mathbf{v}^{(t)}-\mathbf{v}^{(t-1)}\|<\delta,$
where $\delta>0$ is a small predefined tolerance.
The regularization parameter $\varepsilon$ introduces a smoothing effect in the optimal transport problem. As $\varepsilon$ decreases, the solution becomes closer to the exact OT plan, but the algorithm may converge more slowly. Conversely, larger values of $\varepsilon$ result in faster computations, but with a higher approximation error in the transport plan.
4.   Block Coordinate Network Simplex Methods with Basis Variable Succession
The goal of the NS method is to identify the optimal basis variables through iterative pivoting. While each NS iteration efficiently updates the basis structure, computing reduced costs for all variables incurs a cost of $O(N)$ per iteration.
To reduce this per-iteration cost, we adopt the BCD strategy: at each iteration, we restrict attention to a selected subset $\mathcal{H}_{k}\subseteq\mathcal{N}$ of variables and solve a corresponding subproblem using the NS method.
The OT problem is particularly well-suited to such block decomposition. In general LPs with $M$ constraints, computing the reduced cost for a single variable requires $O(M)$ operations. However, due to the special network structure of the OT problem, each reduced cost can be computed in $O(1)$ time via the relation
$r_{ij}=c_{ij}-(\pi_{i}-\pi_{n+j}),$
We now summarize the notation and assumptions used throughout this section:
  • $\mathbf{x}^{k}\in\mathbb{R}^{N}$: the basic feasible solution at iteration $k$,
  • $\mathbf{x}^{*}$:the optimal solution to the full OT problem (2),
  • $\mathcal{B}_{k}\subseteq\{1,\dots,N\}$: the set of basis indices at iteration $k$,
  • $\mathcal{B}^{\ast}\subseteq\{1,\dots,N\}$: the basis of a globally optimal solution,
  • $\mathcal{H}_{k}\subseteq\{1,\dots,N\}$: the block of variable indices selected at iteration $k$,
  • $\bar{c}^{k}:=\bar{c}(\mathbf{x}^{k},\mathcal{B}_{k})$: the reduced cost vector at $\mathbf{x}^{k}$.
4.1 Basis Variables Succession
In BCD schemes for the OT problem, the convergence behaviour hinges on the construction of the working set $\mathcal{H}_{k}$ at each iteration $k$.
The following theorem identifies the precise conditions under which a restricted NS step on a block $\mathcal{H}_{k}$ yields a globally optimal solution. It also shows when global optimality cannot be achieved.
Theorem 2.
Apply the network–simplex method to the subproblem restricted to $\mathcal{H}_{k}$, and set
$d^{k}\in\mathcal{D}\bigl{(}\mathbf{x}^{k};\mathcal{H}_{k}\bigr{)},\qquad\mathbf{x}^{k+1}=\mathbf{x}^{k}+d^{k}.$
Then:
  1. (i)
    If $\mathcal{B}_{k}\subseteq\mathcal{H}_{k}$ and $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$, the iterate $\mathbf{x}^{k+1}$ is globally optimal.
  2. (ii)
    If $\mathcal{B}_{k}\not\subseteq\mathcal{H}_{k}$, then—even when $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$—the iterate $\mathbf{x}^{k+1}$ cannot be globally optimal.
Proof.
(i) Because $\mathcal{B}_{k}\subseteq\mathcal{H}_{k}$, the NS subproblem starts with basis $\mathcal{B}_{k}$. Pivoting within $\mathcal{H}_{k}$ yields a basis ^$\widehat{\mathcal{B}}\subseteq\mathcal{H}_{k}$ satisfying
[…]
Since $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$, the subproblem contains an optimal basis of the full problem; NS therefore stops only when no negative reduced costs remain in $\mathcal{H}_{k}$. As dual variables depend solely on the basis,
^$\widehat{\mathcal{B}}\cap\mathcal{H}_{k}=\mathcal{B}^{\ast},$
so condition (5) extends to all indices $i=1,\dots,N$:
[…]
By complementary slackness, $\mathbf{x}^{k+1}$ is globally optimal.
(ii) Assume $j\in\mathcal{B}_{k}\setminus\mathcal{H}_{k}$ with $\mathbf{x}^{k}_{j}>0$. Because $j\notin\mathcal{H}_{k}$, the subproblem cannot pivot $j$ or alter its value, hence
$\mathbf{x}^{k+1}_{j}=\mathbf{x}^{k}_{j}>0.$
In any optimal OT solution all non–basic variables are zero. Since $j\notin\mathcal{B}^{\ast}$ (because $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$), optimality would require $x^{\ast}_{j}=0$, contradicting $\mathbf{x}^{k+1}_{j}>0$. Therefore $\mathbf{x}^{k+1}$ cannot be optimal. ∎
The above analysis motivates a practical strategy, which we call basis variable succession: at each iteration, we selecte the block $\mathcal{H}_{k}$ includes the basis variables $\mathcal{B}_{k}$ obtained from the previous subproblem. This design contrasts with existing BCD approaches [xie2023randomized] that select blocks randomly.
we can solve the subproblem in iteration $k$ with complexity $O(L_{k}|\mathcal{H}_{k}|)$, which is significantly lower than the $O(LN)$ complexity of solving the full problem directly, where $L_{k}$ and $L$ denote the number of NS iterations for the subproblem and the full problem, respectively.
This strategy offers two practical advantages:
  1. 1.
    The subproblem always includes all constraints and at least $M-1$ variables, ensuring feasibility.
  2. 2.
    Reusing the previous basis variables as the initialization eliminates the need for recomputing a starting basis in the NS method.
4.2 A Prototype of Block Coordinate Descent Network Simplex Methods and its Finite Termination
Following the strategy of §4.1, choose a block
$\mathcal{H}_{k}\supseteq\mathcal{B}_{k},\qquad\lvert\mathcal{H}_{k}\rvert\geq M-1,$
and split it into
$\mathcal{B}_{k},\qquad\mathcal{I}_{k}=\mathcal{H}_{k}\setminus\mathcal{B}_{k}.$
If $\mathcal{I}_{k}$ is picked arbitrarily, the subproblem
$\mathcal{D}\!\bigl{(}\mathbf{x}^{k};\mathcal{H}_{k}\bigr{)}$
might admit only the zero direction, yielding the same basis $\mathcal{B}_{k+1}=\mathcal{B}_{k}$ and $\mathbf{x}^{k}=\mathbf{x}^{k+1}$.
Define the index set of negative reduced costs
$\mathcal{N}_{k}\;=\;\bigl{\{}\,\ell\in\{1,\dots,nm\}:\;\bar{c}^{\,k}_{\ell}<0\bigr{\}}.$
From iteration $k$ on we impose the rule
$\mathcal{I}_{k}\cap\mathcal{N}_{k}\;\neq\;\varnothing.$ (6)
If $\mathcal{N}_{k}=\varnothing$, then $\bar{c}^{\,k}\geq 0$ and, by complementary slackness, $\mathbf{x}^{k}$ is already globally optimal.
Under condition (6) each subproblem produces a strictly improving search direction, ensuring monotone decrease of the objective value. The next theorem establishes finite termination and global optimality of the resulting sequence $\{\mathbf{x}^{k}\}$.
Theorem 3.
The proposed prototype of BCDNS method satisfies the following properties:
  1. 1.
    Monotonicity and Convergence: The objective function decreases monotonically at each iteration, and the method converges to the global optimal solution.
  2. 2.
    Finite Termination: The method terminates within a finite number of iterations.
Proof of Part 1.
Monotonicity: For any iteration $k$, based on the selection rule for $\mathcal{I}_{k}$, there exists at least one $l\in\mathcal{I}_{k}$ such that the reduced cost of $x_{l}$ is negative. By the properties of the reduced cost, adding $x_{l}$ to the basis variables decreases the total cost, i.e.,
$q(\mathbf{x}^{k};\mathcal{B}_{k}\cup\{l\})<q(\mathbf{x}^{k};\mathcal{B}_{k}).$
where $q(\mathbf{x}^{k};\mathcal{I}_{k})$ is defined by (4) as the optimal value of the subproblem associated with the index set $\mathcal{I}_{k}$. Since $l\in\mathcal{I}_{k}$, it follows that:
$q(\mathbf{x}^{k};\mathcal{H}_{k})=q(\mathbf{x}^{k};\mathcal{B}_{k}\cup\mathcal{I}_{k})\leq q(\mathbf{x}^{k};\mathcal{B}_{k}\cup\{l\})<q(\mathbf{x}^{k};\mathcal{B}_{k}).$
Furthermore, since $\mathcal{J}_{k}$ is the optimal basis variables of the previous subproblem, we have:
$q(\mathbf{x}^{k};\mathcal{B}_{k})=q(\mathbf{x}^{k-1};\mathcal{H}_{k-1}).$
Combining the above, it follows that:
$q(\mathbf{x}^{k};\mathcal{H}_{k})<q(\mathbf{x}^{k-1};\mathcal{H}_{k-1}).$
Thus, the cost function decreases monotonically at each iteration.
Global Optimality: After solving each subproblem $\mathcal{D}(\mathbf{x}^{k};\mathcal{H}_{k})$, we compute the reduced costs of all variables. If all reduced costs satisfy $r_{ij}\geq 0$, the algorithm terminates; otherwise, the next iteration begins. By the properties of the reduced cost, this condition guarantees that the final solution is globally optimal. ∎
Proof of Part 2.
The total number of possible combinations of the basis variables among $N$ variables is at most $\binom{N}{M-1}$. Based on the selection rule, in each iteration, the set $\mathcal{I}_{k}$ always contains at least one variable with a negative reduced cost. Denote the variable with the smallest reduced cost as $x_{l}$. When solving the subproblem using the NS method, $x_{l}$ will be added to the basis variables, while one of the current basis variables $\mathcal{B}_{k}$will leave the basis. This implies that each iteration examines at least one new combination of the basis variables. Consequently, after at most $\binom{N}{M-1}$ iterations, all possible combinations of the basis variables will have been examined. Therefore, the algorithm is guaranteed to terminate within a finite number of iterations. ∎
4.3 Grouped Selection Block Coordinate Descent Network Simplex method
Since an optimal solution to the OT problem is typically sparse, most variables take value zero and do not belong to the optimal basis. Consequently, selecting $\mathcal{I}_{k}$ uniformly at random often results in the inclusion of irrelevant variables, slowing convergence.
To address this issue, we propose a grouped block coordinate selection strategy that partitions the variable index set $\{1,\dots,N\}$ into two disjoint subsets:
  • $\mathcal{R}$: indices of variables deemed likely to enter the optimal basis $\mathcal{B}^{*}$,
  • $\mathcal{S}$: indices of variables with low likelihood of being part of the optimal basis $\mathcal{B}^{*}$.
Let $0<s<t\leq 1$ be fixed sampling parameters. At each iteration, the reduced costs are computed for a randomly selected subset of size $tN$ from $\mathcal{R}$, together with a small subset from $\mathcal{S}$ for exploration. Among these, we select the $sN$ variables with the smallest reduced costs to form the block $\mathcal{I}_{k}$, as these variables are most promising for descent directions.
If all selected variables have non-negative reduced costs, the algorithm infers that none are useful for further improvement and moves them to set $\mathcal{S}$. The selection process is then repeated until condition (6) is satisfied.
This strategy guides the search toward structurally relevant directions and improves computational efficiency. Based on this mechanism, we introduce the Grouped Selection Block Coordinate Descent Network Simplex (GS-BCDNS) method. The full algorithm is presented below.
Algorithm 1 Grouped Selection Block Coordinate Descent Network Simplex Method (GS-BCDNS)
Initialization:
Find an initial feasible solution $X^{0}\in\mathbb{R}^{n\times m}$. Set the initial basis variables as $\mathcal{J}_{0}$. Choose parameters $s,t\in(0,1)$ such that $s<t$. Initialize $\mathcal{S}=\emptyset$ and =$\mathcal{R}=\{1,2,\ldots,N\}$. Shuffle $\mathcal{R}$ randomly. Define the threshold sequence $\{e_{k}\}$.
for $k=0,1,2,\ldots$ do
     Step 1: Select the first $\min(|\mathcal{R}|,tN)$ variables from $\mathcal{R}$ and $\min(|\mathcal{S}|,0.1tN)$ variables from $\mathcal{S}$. Compute the reduced costs for all selected variables and sort them in ascending order. Select the $sN$ variables with smallest reduced costs to form $\mathcal{I}_{k}$. If all selected variables have non-negative reduced costs, move them to $\mathcal{S}$ and repeat Step 1.
     Step 2: Move the non-selected variables with positive reduced costs to $\mathcal{S}$, and return variables with non-positive reduced costs to $\mathcal{R}$.
     Step 3: Set $\mathcal{H}_{k}=\mathcal{I}_{k}\cup\mathcal{B}_{k}$.
     Step 4: Use the NS method to solve the subproblem $\mathcal{D}(x^{k};\mathcal{H}_{k})$ and obtain $d^{k}\in\mathcal{D}(x^{k};\mathcal{H}_{k})$. Update the basis to $\mathcal{B}_{k+1}$. Move variables with reduced cost $r_{ij}<e_{k}$ to $\mathcal{R}$, and move the remaining variables to $\mathcal{S}$.
     Step 5: Update $x^{k+1}:=x^{k}+d^{k}$.
     Step 6: Optimality check: If $\mathcal{R}=\emptyset$, compute reduced costs for all variables. If $r_{ij}\geq 0$ for all $i,j$, terminate; otherwise, return to Step 1.
end for
Let $\tilde{k}$ denote the total number of outer iterations, and let $l_{k}$ represent the number of NS iterations required to solve the subproblem at iteration $k$. The time complexity of the GS-BCDNS method is given by:
$O\left(\tilde{k}tN+\sum_{k=1}^{\tilde{k}}l_{k}sN\right).$
The complexity clearly depends on the choice of sampling parameters $s$ and $t$. Decreasing their values reduces the cost per iteration, but increases the total number of iterations $\tilde{k}$. Thus, careful tuning of $s$ and $t$ is essential to balance convergence speed and efficiency.
In each iteration, the space complexity of the GS-BCDNS method is:
$O(|\mathcal{H}_{k}|)=O(sN+M-1).$
Compared to the Sinkhorn algorithm, which requires $O(N)$ space, the GS-BCDNS method is more memory-efficient. However, relative to existing BCD methods, the memory usage depends on the size of the selected block $\mathcal{H}_{k}$ [xie2023randomized].
5.   Numerical Experiments
In this section, we perform numerical experiments on various instances of OT problems1.11footnotetext: 1 Numerical experiments were conducted on a computer with an Intel Core i7-12700H CPU (2.70 GHz) and 16 GB of RAM. The implementation was performed using Python 3.8, and the code has been made publicly available on GitHub at https://github.com/842422492/BCDNS.git. We first solve randomly generated OT problems using the GS-BCDNS method to evaluate the impact of parameters $s$ and $t$ on algorithmic efficiency. Subsequently, we compute the Wasserstein distance between uniform and normal distributions using different algorithms to compare their performance.
5.1 Experiment Settings
For simplicity, we assume that the numbers of supply and demand nodes are equal, i.e., $n=m$.
Test Problem 1: Ten randomly generated discrete OT problems using randomly sampled numbers. $n=1000$.
Test Problem 2: The Wasserstein distance between a one-dimensional uniform distribution on $[-1,1]$ and a standard normal distribution on $\mathbb{R}$. $n=1000$.
Test Problem 3: The Wasserstein distance between a two-dimensional uniform distribution on $[-1,1]^{2}$ and a standard normal distribution on $\mathbb{R}^{2}$. $n=1000$.
Test Problem 4: The Wasserstein distance between a three-dimensional uniform distribution on $[-1,1]^{3}$ and a standard normal distribution on $\mathbb{R}^{3}$. $n=1000$.
Test Problem 5: The Wasserstein distance between a one-dimensional uniform distribution on $[-1,1]$ and a standard normal distribution for different problem sizes. $n$ starts from 50 and increases by 50 up to 300.
The discrete OT problems in Test Problem 1 were generated as follows: two sets of $n=1000$ random numbers were sampled from a uniform distribution on $[1,1000]$ to represent the supply and demand distributions $P$ and $Q$, respectively. Then adjuste the two sets to ensure that their sums were equal. Additionally, a cost matrix of size $1000\times 1000$ was generated from the same uniform distribution.
The Wasserstein distances in Test Problem 2, 3, and 4 were estimated as follows: $n$ samples were randomly drawn from the uniform and normal distributions, denoted as $y$ and $\tilde{y}$, respectively. For each pair, i.i.d. samples $\{\tilde{y}_{i}\}$ and $\{y_{j}\}$, $1\leq i,j\leq 1000$, were used to approximate the two continuous probability measures. The cost matrix was defined as: $c_{i,j}=\text{dist}(\tilde{y}_{i},y_{j})^{2},$ and the probability mass vectors were set as $\mathbf{p}=\mathbf{q}=(1/n,\dots,1/n)^{T}\in\mathbb{R}^{n}$.
Figure 1: The point distributions in Test Problem 2, 3 and 4
Algorithm Settings:
  • Sinkhorn (Stabilized) Algorithm: As used in [xie2023randomized], we follow the method proposed in [cuturi2013sinkhorn] for computing the Wasserstein distance. Let $\gamma$ be the entropy regularization coefficient, which is set to $\gamma=1/(4\log n)$ as suggested in [dvurechensky2018computational]. We consider the parameter settings $\varepsilon=10^{-4},10^{-3},0.01,0.1$. Sinkhorn iterations are projected onto the feasible region using a rounding procedure, specifically Algorithm 2 in [altschuler2017near]. To mitigate numerical instability, all updates are performed in log space using the LogSumExp function. The algorithm terminates when either the change in the objective function value falls below $\delta=10^{-6}$ or the runtime exceeds 2000 seconds.
  • GS-BCDNS: When solving Test Problem 1, the parameter $s$ was varied from $1/n$ to 0.9, while the parameter $t$ was varied from $2/n$ to 1. For Test Problems 2, 3, 4 and 5, the values of $s$ and $t$ were fixed at $5/n$ and $10/n$, respectively. For each Test Problem, set $\{e_{k}\}$ to positive infinity and calculate the initial basic feasible solution using the Northwest Corner Rule.
5.2 Effect of Parameters $s$ and $t$
In the BCDNS method, the parameters $s$ and $t$ play a crucial role. Here, we vary the parameters $s$ and $t$ to solve Test Problem 1. The number of reduced cost computations serves as a key indicator in the BCDNS method. For each case, we record the corresponding number of pivots and reduced cost computations. And then we calculate the averages of these metrics to evaluate the algorithm’s performance. In this context, the number of pivots refers to the number of iterations performed by the NS method.
Figure 2: Effects of $s$ and $t$ on the number of pivots and reduced cost computations
Figure 1 illustrates the effects of $s$ and $t$ on two key performance metrics: the number of pivots and the total number of reduced cost computations. We can observe that the number of pivots decreases as $s$ increases, as larger subproblem sizes reduce the total iterations required by the NS method. However, as $s$ increases, the size of each subproblem also grows, which, while reducing the number of pivots, increases the computational cost of each pivot. Similarly, increasing $t$ enhances the quality of the variables selected for the subproblem, further reducing the number of pivots.
For small values of $s$, increasing $s$ reduces the total number of reduced cost computations by enabling more efficient solutions to larger subproblems. However, beyond a certain threshold, further increases in $s$ lead to an overall rise in computational effort due to the greater complexity of individual subproblems. The effect of $t$ is less pronounced, except when $s$ is very small, in which case increasing $t$ significantly increases the computational burden, as the algorithm approaches the behavior of the standard NS method. Extensive numerical experiments consistently reveal that reduced cost computations are minimized when $t$ is approximately $2s$.
5.3 Comparison with the Existing Methods
In this chapter, we primarily evaluate the performance differences among the GS-BCDNS method, the Sinkhorn algorithm, and the NS method.
Figure 3: Comparison of algorithms to computer Wasserstein distance
As described in Section 4.1, we apply the Sinkhorn algorithm with different values of $\epsilon$ and the GS-BCDNS method to solve Test Problems 2, 3 and 4. The stopping criterion for the Sinkhorn algorithm is set such that the algorithm terminates when the change in the objective function value falls below $\delta=10^{-6}$ or when the runtime exceeds 2000 seconds. The GS-BCDNS method is tested with fixed parameters $s=5/n$ and $t=10/n$. In the evaluation, the x-axis represents the runtime, while the y-axis represents the gap between the current objective function value and the optimal objective function value, where the optimal solution is obtained using the interior-point method implemented in Python’s SciPy library.
The results in Figure 3 indicate that for larger values of $\epsilon$ (i.e, $\epsilon=0.1$ and $\epsilon=0.01$), the Sinkhorn algorithm exhibits a faster convergence speed than the GS-BCDNS method; however, the approximation error remains relatively large. When $\epsilon=0.001$, the convergence speed of the Sinkhorn algorithm is comparable to that of the GS-BCDNS method, but its final convergence is slightly slower. For extremely small values of $\epsilon$, the convergence speed of the Sinkhorn algorithm is significantly slower than that of the GS-BCDNS method.
Figure 4: Comparison of algorithms to solve Test Problem 5
Figure 4 presents the results of solving Test Problem 5 using both the NS method and the GS-BCDNS method. The y-axis of Figure 4 represents the ratio of the NS method’s runtime to that of the GS-BCDNS method. The results demonstrate that the GS-BCDNS method significantly outperforms the NS method across all tested datasets.
6.   Conclusion
In this paper, we review the existing methods for solving the OT problems, including the NS method, the BCD method, and the Sinkhorn algorithm. Then we propose the BCDNS methods and explain thier theoretical foundations and the basis variable succession strategy. Furthermore, we provide a detailed algorithmic framework and established theoretical guarantees for finite termination and convergence to the exact optimal solution.
Extensive numerical experiments demonstrate that the BCDNS method achieves superior performance in terms of runtime and computational cost compared to the NS method. While the BCDNS method converges slightly slower than the Sinkhorn algorithm with large values of $\epsilon$, it has the advantage of guaranteeing convergence to an exact solution. Moreover, compared to the Sinkhorn algorithm with small values of $\epsilon$, the BCDNS method exhibits even faster convergence. This makes the BCDNS method particularly valuable for large-scale OT problems that require exact solutions.
Future work could explore several directions. First, improving the selection strategy for block coordinates may further enhance convergence speed. Second, extending the BCDNS method to more general network flow problems could broaden its applicability.
Appendix A Proof of $\operatorname{rank}(A)=M-1$
Proof.
Let
$A=\begin{bmatrix}I_{n}&I_{n}&\cdots&I_{n}\\ -\mathbf{1}_{n}^{\top}&0&\cdots&0\\ 0&-\mathbf{1}_{n}^{\top}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&-\mathbf{1}_{n}^{\top}\end{bmatrix}\in\mathbb{R}^{M\times N},\quad M=n+m,\;N=nm,$
where $I_{n}$ is the $n\times n$ identity matrix and $\mathbf{1}_{n}$ is the $n$-vector of all ones.
(i) $\operatorname{rank}(A)\leq M-1$.
Denote the rows of $A$ by $r_{1},\dots,r_{M}$. Observe that
[…]
because each of the first $n$ rows contributes exactly one “$+1$” in each block and each of the last $m$ rows contributes exactly one “$-1$”, so they cancel out. Hence the rows are linearly dependent, and
$\operatorname{rank}(A)\;\leq\;M-1.$
(ii) $\operatorname{rank}(A)\geq M-1$.
Interpret $A$ as the node–arc incidence matrix of the complete bipartite directed graph $K_{n,m}$ with $M=n+m$ nodes and $nm$ arcs. Choose any spanning tree $T$ of this graph, which has exactly $M-1$ edges. Let
$B\;=\;A(:,E(T))\;\in\;\mathbb{R}^{M\times(M-1)}$
be the submatrix of $A$ whose columns correspond to the edges in $T$. Since $T$ is a tree, its incidence matrix has rank $M-1$, and in particular there is an $(M-1)\times(M-1)$ submatrix of $B$ with determinant ±$\pm 1$. Therefore
$\operatorname{rank}(A)\;\geq\;M-1.$
Combining (i) and (ii) gives
$\operatorname{rank}(A)\;=\;M-1.$
[…]

### v3

@@
-references.bib
+Block Coordinate Descent Network Simplex Methods for Optimal Transport
 Abstract
-We propose the Block Coordinate Descent Network Simplex (BCDNS) method for solving large-scale discrete Optimal Transport (OT) problems.
+We propose the Block Coordinate Descent Network Simplex (BCDNS) method for solving large-scale Optimal Transport (OT) problems.
 BCDNS integrates the Network Simplex (NS) algorithm with a block coordinate descent (BCD) strategy, decomposing the full problem into smaller subproblems per iteration and reusing basis variables to ensure feasibility.
 We prove that BCDNS terminates in a finite number of iterations with an exact optimal solution, and we characterize its per-iteration complexity as $O(sN)$, where $s\in(0,1)$ is a user-defined parameter and $N$ is the total number of variables.
-Numerical experiments demonstrate that BCDNS matches the classical NS method in solution accuracy, reduces memory footprint compared to Sinkhorn algorithm, achieves speed-ups of up to tens of times over the classical NS method, and exhibits runtime comparable to high-precision Sinkhorn algorithm.
+Numerical experiments show that BCDNS attains the same exact objective value as the classical NS method while substantially reducing runtime on large instances, with speed-ups of up to tens of times over NS.
+In addition, comparisons with the stabilized Sinkhorn algorithm under different regularization levels illustrate that BCDNS converges to the unregularized OT optimum, and its runtime is competitive with Sinkhorn when high accuracy is required.
 Keywords: Optimal Transport, Network Simplex, Block Coordinate Descent, Sinkhorn, Wasserstein Distance
@@
 It is rooted in mathematical optimization and has a rich history dating back to the 18th century.
-Gaspard Monge [monge1781deblais] first proposed it in 1781.
-Then in the mid-20th century, Leonid Kantorovich [kantorovich1942translocation] extended Monge’s framework to encompass continuous distributions, reformulating it as a linear programming model and significantly broadening its theoretical foundation and practical applications.
-In this thesis, we focus specifically on solving the discrete OT problem.
+Gaspard Monge [13] first proposed it in 1781.
+Then in the mid-20th century, Leonid Kantorovich [10] extended Monge’s framework to encompass continuous distributions, reformulating it as a linear programming model and significantly broadening its theoretical foundation and practical applications.
+In this paper, we focus specifically on solving the discrete OT problem.
 The OT problem has found applications in various disciplines.
-In machine learning, the OT problem has been used for tasks such as domain adaptation [courty2016optimal] and generative modeling  [arjovsky2017wasserstein].
-In economics, the OT problem has provided insights into resource allocation, market equilibrium, and pricing mechanisms [galichon2016optimal].
-In biology, the OT problem has been applied to analyze genetic expression patterns [schiebinger2019optimal].
-The OT problem has found applications in various disciplines.
-In machine learning, the OT problem has been used for tasks such as domain adaptation [courty2016optimal] and generative modeling [arjovsky2017wasserstein].
-In economics, the OT problem has provided insights into resource allocation, market equilibrium, and pricing mechanisms [galichon2016optimal].
-In biology, the OT problem has been applied to analyze genetic expression patterns [schiebinger2019optimal].
-In geophysics and signal processing, the OT problem has been employed for seismic waveform comparison and inversion.
-[liao2022fast].
+In machine learning, the OT problem has been used for tasks such as domain adaptation [5] and generative modeling [3].
+In economics, the OT problem has provided insights into resource allocation, market equilibrium, and pricing mechanisms [8].
+In biology, the OT problem has been applied to analyze genetic expression patterns [14].
+In geophysics and signal processing, the OT problem has been used for seismic waveform comparison and inversion [12].
 Several methods have been developed to solve the OT problem.
-Each method has distinct strengths and limitations.
-The Hungarian method is a classical approach for solving linear assignment problems and can also be applied to OT problems [kuhn1955hungarian].
-It achieves exact optimal solutions by constructing a cost matrix and iteratively reducing reducing dimensions of the matrix.
-However, its time complexity is $O(N^{2})$, where $N$ is the number of variables in OT problem, and hence it is not applicable for the large-scale OT problem.
-The Sinkhorn algorithm is a popular method for the large-scale OT problem due to its computational efficiency [cuturi2013sinkhorn].
+Each method has different strengths and limitations.
+The Hungarian method is a classical algorithm for the linear assignment problem [11], which can be viewed as a special case of OT where the transport plan is restricted to a one-to-one matching.
+It produces an exact optimum in $O(n^{3})$ time for an $n\times n$ assignment instance.
+However, for general OT, the Hungarian method is no longer applicable without expanding the problem size, which becomes prohibitive for large-scale problems.
+The Sinkhorn algorithm is a popular method for the large-scale OT problem due to its computational efficiency [6].
 It solves the OT problems with regularization of entropy, which is an approximation of the OT problem.
 Therefore it provides approximate solutions of the OT problem, and it may not be suitable for applications that require exact solutions.
-In recent years, many improvements based on the Sinkhorn algorithm have been proposed, such as the Greenkhorn algorithm [altschuler2017Near] and the Adaptive Primal–Dual Accelerated Gradient Descent (APDAGD) algorithm [dvurechensky2018Computational].
+In recent years, many improvements based on the Sinkhorn algorithm have been proposed, such as the Greenkhorn algorithm [2] and the Adaptive Primal–Dual Accelerated Gradient Descent algorithm [7].
 However, these methods still operate within the Sinkhorn framework and do not overcome the limitation of producing only approximate OT solutions.
-In contrast, the Network Simplex (NS) method, a specialized variant of the simplex method, is well known to yield exact solutions.
-Michael D.
-Grigoriadis [grigoriadis1986efficient] reported that the NS method is reliable among several algorithms.
-However, the NS method is not a polynomial time algorithm and it takes $O(N)$ time in each iteration, which leads to slow computational performance when applied to large-scale problems.
-We note that the above three methods needs at least $O(N)$ space complexity.
-To overcome the difficulty of large space complexity for the OT problem, Yue Xie et al.
-[xie2023randomized] proposed the Block Coordinate Descent (BCD) method for solving the OT problem.
-Their method constructs a small subproblem by randomly selecting variables based on predefined rules and solves it in each iteration.
-Since the subproblem is small, the BCD method does not require the O(N) space complexity.
+In contrast, the Network Simplex (NS) method [9], a specialized variant of the simplex method, is well known to yield exact solutions.
+However, the NS method takes $O(N)$ time in each iteration, which leads to slow computational performance when applied to large-scale problems.
+Yue Xie et al.
+[18] proposed a block coordinate descent (BCD) approach for solving large-scale OT problems.
+At each iteration, the method constructs a restricted subproblem by selecting a subset of variables according to a prescribed sampling rule and then solves the subproblem while keeping the remaining variables fixed.
+By operating on a much smaller working set, the BCD approach can substantially reduce the computational burden per-iteration and, in practice, also reduces working memory compared to methods that process all $N$ variables at once.
 However, since the BCD method proposed by Yue Xie et al.
@@
 As data sizes continue to grow, we recognize the increasing importance of solving large-scale OT problems with both speed and accuracy.
-To address these challenges, we aim to develop a method that enables fast computation of exact solutions for large-scale OT problems.
+To address these challenges, our aim is to develop a method that enables fast computation of exact solutions for large-scale OT problems.
 To achieve the goal of efficiently solving large-scale OT problems, we introduce the idea of the BCD method.
-The BCD method divides a large-scale OT problem into smaller subproblems, enabling a reduction in computational complexity.
-Although the BCD method is generally considered slow when exact solutions are required, we observe that the optimal solutions of the OT problem are typically sparse—due to the relatively small number of equality constraints compared to the number of variables—which significantly improves the efficiency of BCD.
+The BCD method divides a large-scale OT problem into smaller subproblems, allowing a reduction in computational complexity.
+Although the BCD method is generally considered slow when exact solutions are required, we observe that the optimal solutions of the OT problem are typically sparse, which significantly improves the efficiency of the BCD method.
 By employing the NS method to solve the subproblems, we guarantee that the final solution is also exact.
-In this paper, we propose the Block Coordinate Descent Network Simplex (BCDNS) method, a novel algorithm that integrates the BCD and NS methods to achieve computational efficiency and solution accuracy.
-Our approach reduces computational complexity by decomposing large-scale OT problems into smaller subproblems, while the NS method ensures that each subproblem is solved with precision.
-Numerical experiments demonstrate significant performance gains of the BCDNS method over traditional approaches, addressing key limitations of existing methods.
-In the BCD method, the selection of the block coordinate is crucial.
-Compared to existing BCD method [xie2023randomized] that randomly selects variables based on matrix properties, we propose that the BCDNS method first selects the basis variables from the previous subproblem as the initial solution for the next subproblem.
-This strategy ensures that the subproblem always has a feasible and an optimal solution.
-Moreover, by integrating the principles of the NS method, we ensure that the BCDNS method converges to the exact optimal solution within a finite number of computations.
-Furthermore, we categorize variables into two groups based on their likelihood of becoming optimal basis variables.
-This targeted selection process prioritizes variables with a higher probability of being optimal basis variables, thereby improving the efficiency of finding the optimal solution.
-The outline of this paper is as follows.
-In Section 2, we introduce the foundational concepts of the OT problem and review several existing methods for solving it.
-In Section 3, we describe the core principles of the proposed BCDNS method.
-We first present its prototype along with the corresponding theoretical proofs.
-Then we develop a concrete implementation based on the prototype.
-Section 4 reports numerical experiments, including parameter estimation for the BCDNS method and a comparative performance analysis against existing approaches.
-Finally, Section 5 concludes the paper by summarizing the key contributions and discussing potential directions for future research.
+In this paper, we develop the Block Coordinate Descent Network Simplex (BCDNS) method for solving large-scale OT problems with exact optimality.
+Compared to existing BCD method [18] that randomly selects variables based on matrix properties, we introduced a basis variable succession strategy to reuse basis information across iterations, which ensures monotone improvement and finite termination.
+Furthermore, we propose a grouped variable selection mechanism to reduce unnecessary reduced-cost evaluations.
+Theoretical analysis establishes convergence to the global optimum in finitely many iterations, and numerical experiments demonstrate that the proposed method significantly outperforms the classical NS method and is competitive with high-precision Sinkhorn algorithms in terms of runtime when high accuracy is required.
+The remainder of this paper is organized as follows.
+Section 2 formulates the discrete OT problem and introduces the notation used throughout the paper.
+Section 3 reviews representative approaches for OT, including the NS method, the BCD methods, and the Sinkhorn algorithm.
+Section 4 presents the proposed BCDNS framework, together with concrete block selection strategies and theoretical properties.
+Section 5 reports the numerical experiments, including the test settings and the empirical results.
+Finally, Section 6 concludes the paper and discusses future directions.
 2.  Discrete Optimal Transport
-We consider a discrete OT problem [villani2009optimal] between $n$ supply locations and $m$ demand locations.
+We consider a discrete OT problem [17] between $n$ supply locations and $m$ demand locations.
 Let $\{1,2,\dots,n\}$ denote the index set of supply points, and $\{1,2,\dots,m\}$ denote that of demand points.
@@
     $A=\begin{bmatrix}I_{n}&I_{n}&\cdots&I_{n}\\ -1_{n}^{\top}&0&\cdots&0\\ 0&-1_{n}^{\top}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&-1_{n}^{\top}\end{bmatrix}\in\mathbb{R}^{M\times N},$
-    where $I_{n}$ denotes the $n\times n$ identity matrix, and $1_{n}$ is an $n$-dimensional column vector of ones.
-With this notation, we can rewrite the OT problem as:
+    where $I_{n}$ denotes the identity matrix $n\times n$, and $1_{n}$ is a column vector of dimensions $n$.
+With this notation, we can rewrite the OT problem as follows:
 $\min_{\mathbf{x}\in\mathbb{R}^{N}}\quad$ $\mathbf{c}^{\top}\mathbf{x}$ (2)
@@
 In this section, we review several representative algorithms for solving the discrete OT problem.
-These include the classical NS method, the Sinkhorn algorithm based on entropy regularization which is widely used for large-scale problems, and the recently proposed BCD method.
+These include the classical NS method, the Sinkhorn algorithm based on entropy regularization, which is widely used for large-scale problems, and the recently proposed BCD method.
 3.1 Network Simplex Method for Optimal Transport
-As shown in Appendix A, the constraint matrix $A\in\mathbb{R}^{M\times N}$ in the OT problem has rank $M-1$.
-This allows us to partition the set of variable indices $\mathcal{N}=\{1,\dots,N\}$ into two subsets: a basis index set $\mathcal{B}\subseteq\mathcal{N}$ with $|\mathcal{B}|=M-1$, and its complement $\mathcal{N}\setminus\mathcal{B}$.
+It is well known that the node–arc incidence matrix of a connected graph has rank $\lvert V\rvert-1$ [1].
+Since the OT network is the complete bipartite graph with $M=n+m$ nodes, we have $\operatorname{rank}(A)=M-1$.
+This allows us to partition the set of variable indices $\mathcal{N}=\{1,\dots,N\}$ into two subsets: a set of basis indices $\mathcal{B}\subseteq\mathcal{N}$ with $|\mathcal{B}|=M-1$, and its complement $\mathcal{N}\setminus\mathcal{B}$.
 The columns of $A$ corresponding to $\mathcal{B}$, denoted $A_{B}\in\mathbb{R}^{M\times(M-1)}$, form a full-rank submatrix.
-The basic variables $\mathbf{x}_{B}$ can then be uniquely determined by the system $A_{B}\mathbf{x}_{B}=\mathbf{b}$, while the non-basic variables $\mathbf{x}_{N}$ are fixed at zero.
-A basic feasible solution further satisfies the non-negativity constraint $\mathbf{x}_{B}\geq 0$.
-To determine whether a given basic solution is optimal, we examine its reduced-cost vector.
-Define the dual potential vector $\boldsymbol{\pi}\in\mathbb{R}^{M}$ via the dual feasibility condition $\boldsymbol{\pi}=(A_{B}^{\top})^{-1}\mathbf{c}_{B}$, where $\mathbf{c}_{B}$ is the cost subvector corresponding to the basis $B$.
-Then, for any index $k\in N\setminus B$, the reduced cost is computed as $r_{k}=c_{k}-\boldsymbol{\pi}^{\top}A_{k}$.
-Due to the structure of the OT problem, each column $A_{ij}$ has exactly one $+1$ and one $-1$, so the reduced cost simplifies to the closed-form expression
+The basic variables $\mathbf{x}_{B}$ are determined by the equality constraints
+$A_{B}\mathbf{x}_{B}=\mathbf{b},$
+while the nonbasic variables satisfy $\mathbf{x}_{N}=\mathbf{0}$.
+A basic feasible solution (BFS) additionally requires $\mathbf{x}_{B}\geq\mathbf{0}$.
+To assess the optimality of a BFS, we examine the reduced costs derived from dual variables(node potentials).
+Let $\boldsymbol{\pi}\in\mathbb{R}^{M}$ denote a vector of node potentials.
+For a given basis $B$ (which corresponds to a spanning tree), we choose $\boldsymbol{\pi}$ to satisfy the dual optimality equations
+$A_{B}^{\top}\boldsymbol{\pi}=\mathbf{c}_{B}.$ (3)
+Since $A_{B}$ has rank M1M-1, the solution of (3) is unique up to an additive constant; we fix this degree of freedom by setting the potential of an arbitrary reference node to zero.
+In practice, $\boldsymbol{\pi}$ can be computed in $O(M)$ time by traversing the spanning tree.
+Given $\boldsymbol{\pi}$, the reduced cost of any variable $\ell$ is
+$\bar{c}_{\ell}\;=\;c_{\ell}-A_{\ell}^{\top}\boldsymbol{\pi},$ (4)
+where $A_{\ell}$ denotes the $\ell$-th column of $A$.
+Equivalently, in vector form $\bar{\mathbf{c}}=\mathbf{c}-A^{\top}\boldsymbol{\pi}$.
+Due to the structure of the OT problem, each column $A_{ij}$ has exactly one +1+1 and one 1-1, so the reduced cost simplifies to the closed-form expression
 $r_{ij}=c_{ij}-(\pi_{i}-\pi_{n+j}).$
-Optimality of the current basic solution is certified if all reduced costs are non-negative, i.e., $r_{ij}\geq 0$ for all $i,j$.
-If a variable $x_{i^{*}j^{*}}$ has $r_{i^{*}j^{*}}<0$, it indicates a direction of descent in the objective, and that variable is introduced into the basis.
+A BFS is optimal if and only if $\bar{c}_{\ell}\geq 0$ for all nonbasic indices $\ell\in N$, i.e., $r_{ij}\geq 0$ for all $i,j$.
 To perform a pivot, we add the index $(i^{*},j^{*})$ to the basis $B$, inducing a unique cycle in the associated flow graph.
-Then determine the step size $\Delta$ by the minimum flow along the negative edges of the cycle, ensuring feasibility.
-Subsequently, update the flow: $x_{i^{*}j^{*}}\leftarrow\Delta$, and for each edge $(u,v)$ along the cycle, the value of $x_{uv}$ is incremented or decremented by $\Delta$ depending on the cycle direction.
-Finally, the variable that reaches zero and remove it from the basis, completing the pivot step [vanderbei2001linear].
+Then choose the step size $\Delta$ as the largest admissible increase along the fundamental cycle, i.e.,
+$\Delta\;=\;\min\{\,x_{e}:e\in\mathcal{C}^{-}\,\},$
+where $\mathcal{C}^{-}$ denotes the set of cycle edges on which the flow is decreased.
+This guarantees that the updated flow remains feasible and nonnegative.
+Subsequently, update the flow: $x_{i^{*}j^{*}}\leftarrow\Delta$, and for each edge $(u,v)$ along the cycle, the value of $x_{uv}$ increases or decreases by $\Delta$ depending on the direction of the cycle.
+Finally, the variable that reaches zero and removes it from the basis, completing the pivot step [16].
 Each iteration of the NS method involves computing the dual potential vector $\boldsymbol{\pi}$ in $O(M)$ time, evaluating all reduced costs in $O(N)$, and performing pivoting operations in $O(M)$.
-Since each pivot strictly improves the objective function by at least $-r_{i^{*}j^{*}}>0$, and the number of possible basic feasible solutions is finite, the NS method terminates in a finite number of iterations.
 3.2 Block Coordinate Descent Method for Optimal Transport
-In this subsection we introduce the BCD method for the OT problem in the LP form (2).
+In this subsection, we introduce the BCD method for the OT problem in the LP form (2).
 The BCD method iteratively selects a subset of coordinates $\mathcal{H}\subseteq\mathcal{N}=\{1,2,\dots,N\}$. For any feasible solution $\mathbf{x}$, we define a subproblem of the BCD method
-$\𝒩D(\mathbf{x};\mathcal{H})\triangleq\arg\min_{\mathbf{d}\in\mathbb{R}^{N}}\left\{\mathbf{c}^{T}(\mathbf{x}+\mathbf{d})\mid\mathbf{x}+\mathbf{d}\geq 0,\;A\mathbf{d}=0,\;d_{i}=0,\;\forall i\in\mathcal{N}\backslash\mathcal{H}\right\}.$ (3)
-$\𝒩q(\mathbf{x};\mathcal{H})\triangleq\min_{\mathbf{d}\in\mathbb{R}^{N}}\left\{\mathbf{c}^{T}(\mathbf{x}+\mathbf{d})\mid\mathbf{x}+\mathbf{d}\geq 0,\;A\mathbf{d}=0,\;d_{i}=0,\;\forall i\in\mathcal{N}\backslash\mathcal{H}\right\}.$ (4)
-Here, $D(\mathbf{x};\mathcal{H})$ represents the set of optimal solutions for the subproblem with $\mathcal{H}$, and $q(\mathbf{x};\mathcal{H})$ denotes the optimal objective value corresponding to $d\in D(\mathbf{x};\mathcal{H})$ [tseng2009block].
+$D(\mathbf{x};\mathcal{H})\triangleq\arg\min_{\mathbf{d}\in\mathbb{R}^{N}}\left\{\mathbf{c}^{T}(\mathbf{x}+\mathbf{d})\mid\mathbf{x}+\mathbf{d}\geq 0,\;A\mathbf{d}=0,\;d_{i}=0,\;\forall i\in\mathcal{N}\backslash\mathcal{H}\right\}$.
+(5)
+$q(\mathbf{x};\mathcal{H})\triangleq\min_{\mathbf{d}\in\mathbb{R}^{N}}\left\{\mathbf{c}^{T}(\mathbf{x}+\mathbf{d})\mid\mathbf{x}+\mathbf{d}\geq 0,\;A\mathbf{d}=0,\;d_{i}=0,\;\forall i\in\mathcal{N}\backslash\mathcal{H}\right\}$.
+(6)
+Here, $D(\mathbf{x};\mathcal{H})$ represents the set of optimal solutions for the subproblem with $\mathcal{H}$, and $q(\mathbf{x};\mathcal{H})$ denotes the optimal objective value corresponding to $d\in D(\mathbf{x};\mathcal{H})$ [15].
 The iterative BCD algorithm proceeds as follows:
@@
   • Optimality Check: If optimality is achieved, terminate the algorithm. Else adjust the block $\mathcal{H}_{k}$ for the next iteration using predefined rules or adaptive heuristics.
-The existing BCD method [xie2023randomized], randomly selects block coordinates based on certain properties of specific matrices.
-In [xie2023randomized], the authors propose the Accelerated Random Block Coordinate Descent algorithm and establish the following convergence theorem.
+The existing BCD method [18], randomly selects block coordinates based on certain properties of specific matrices.
+In [18], the authors propose the Accelerated Random Block Coordinate Descent algorithm and establish the following convergence theorem.
 Theorem 1.
-The gap between the current objective function value and the optimal objective function value converges with rate $1-v$, i.e.,
+The gap between the current objective function value and the optimal objective function value converges with rate 1v1-v, i.e.,
 $\mathbf{c}^{\top}\mathbf{x}^{k+1}-\mathbf{c}^{\top}\mathbf{x}^{*}\leq(1-v)(\mathbf{c}^{\top}\mathbf{x}^{k}-\mathbf{c}^{\top}\mathbf{x}^{*}).$
@@
 These approaches provide a simple mechanism for constructing subproblems by selecting block coordinates.
-However, when the size n of the OT problems is large, $v$ becomes extremely small, which leads to slow convergence.
+However, when the size n of the OT problems is large, $v$ becomes extremely small, leading to slow convergence.
 Furthermore, these approaches do not account for the sparsity of the solution of the OT problem.
 Many variables that ultimately take zero values continue to be included in the block coordinates at each iteration.
-Moreover, these approaches have an inherent limitation: they cannot guarantee convergence to the optimal solution within a finite number of iterations.
-This limitation arises from the possibility of extreme cases where the randomly selected block of coordinates rare includes the optimal basis variables.
+Moreover, these approaches typically provide only asymptotic convergence guarantees, rather than finite termination at an exact optimal solution.
 3.3 Sinkhorn Algorithm for Optimal Transport
-Instead of solving the classical OT problem defined in (1), the Sinkhorn algorithm [cuturi2013sinkhorn] introduces an entropy regularization term
+Instead of solving the classical OT problem defined in (1), the Sinkhorn algorithm [6] introduces an entropy regularization term
 $\min_{X\in\mathbb{R}_{+}^{n\times m}}\;\sum_{i=1}^{n}\sum_{j=1}^{m}c_{ij}x_{ij}-\varepsilon H(X),$
@@
 where $\delta>0$ is a small predefined tolerance.
-The regularization parameter $\varepsilon$ introduces a smoothing effect in the optimal transport problem.
+The regularization parameter $\varepsilon$ introduces a smoothing effect in the OT problem.
 As $\varepsilon$ decreases, the solution becomes closer to the exact OT plan, but the algorithm may converge more slowly.
@@
   • $\mathbf{x}^{k}\in\mathbb{R}^{N}$: the basic feasible solution at iteration $k$,
-  • $\mathbf{x}^{*}$:the optimal solution to the full OT problem (2),
+  • $\mathbf{x}^{*}$:the optimal solution to the OT problem (2),
   • $\mathcal{B}_{k}\subseteq\{1,\dots,N\}$: the set of basis indices at iteration $k$,
@@
 In BCD schemes for the OT problem, the convergence behaviour hinges on the construction of the working set $\mathcal{H}_{k}$ at each iteration $k$.
-The following theorem identifies the precise conditions under which a restricted NS step on a block $\mathcal{H}_{k}$ yields a globally optimal solution.
-It also shows when global optimality cannot be achieved.
+The following theorem clarifies when solving a subproblem with the NS method restricted to $\mathcal{H}_{k}$ can yield a globally optimal solution of the full OT problem, and when this is impossible.
 Theorem 2.
-Apply the network–simplex method to the subproblem restricted to $\mathcal{H}_{k}$, and set
+Apply the NS method to the subproblem restricted to $\mathcal{H}_{k}$, and set
 $d^{k}\in\mathcal{D}\bigl(\mathbf{x}^{k};\mathcal{H}_{k}\bigr),\qquad\mathbf{x}^{k+1}=\mathbf{x}^{k}+d^{k}.$
@@
   1. (i)
-    If $\mathcal{B}_{k}\subseteq\mathcal{H}_{k}$ and $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$, the iterate $\mathbf{x}^{k+1}$ is globally optimal.
+    If $\mathcal{B}_{k}\subseteq\mathcal{H}_{k}$ and $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$, then $\mathbf{x}^{k+1}$ is a globally optimal solution of the full OT problem.
   2. (ii)
-    If $\mathcal{B}_{k}\not\subseteq\mathcal{H}_{k}$, then—even when $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$—the iterate $\mathbf{x}^{k+1}$ cannot be globally optimal.
+    Assume that the global optimal solution $\mathbf{x}^{\ast}$ is unique.
+If $\mathcal{B}_{k}\not\subseteq\mathcal{H}_{k}$, then even when $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$, the iterate $\mathbf{x}^{k+1}$ cannot be globally optimal.
 Proof.
-(i) Because $\mathcal{B}_{k}\subseteq\mathcal{H}_{k}$, the NS subproblem starts with basis $\mathcal{B}_{k}$.
-Pivoting within $\mathcal{H}_{k}$ yields a basis ^$\widehat{\mathcal{B}}\subseteq\mathcal{H}_{k}$ satisfying
-[…]
-Since $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$, the subproblem contains an optimal basis of the full problem; NS therefore stops only when no negative reduced costs remain in $\mathcal{H}_{k}$.
-As dual variables depend solely on the basis,
-^$\widehat{\mathcal{B}}\cap\mathcal{H}_{k}=\mathcal{B}^{\ast},$
-so condition (5) extends to all indices $i=1,\dots,N$:
-[…]
-By complementary slackness, $\mathbf{x}^{k+1}$ is globally optimal.
-(ii) Assume $j\in\mathcal{B}_{k}\setminus\mathcal{H}_{k}$ with $\mathbf{x}^{k}_{j}>0$. Because $j\notin\mathcal{H}_{k}$, the subproblem cannot pivot $j$ or alter its value, hence
-$\mathbf{x}^{k+1}_{j}=\mathbf{x}^{k}_{j}>0.$
-In any optimal OT solution all non–basic variables are zero.
-Since $j\notin\mathcal{B}^{\ast}$ (because $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$), optimality would require $x^{\ast}_{j}=0$, contradicting $\mathbf{x}^{k+1}_{j}>0$.
-Therefore $\mathbf{x}^{k+1}$ cannot be optimal.
+(i) Because $\mathcal{B}_{k}\subseteq\mathcal{H}_{k}$, the current basic feasible solution $\mathbf{x}^{k}$ is feasible for the restricted subproblem, and thus the NS method can be initialized from $\mathbf{x}^{k}$ without losing feasibility.
+Let $\mathbf{x}^{k+1}$ be an optimal solution of the restricted subproblem.
+Then $\mathbf{x}^{k+1}$ is feasible for the full OT problem (it satisfies $A\mathbf{x}^{k+1}=b$ and $\mathbf{x}^{k+1}\geq 0$), hence
+$c^{\top}\mathbf{x}^{k+1}\ \geq\ c^{\top}\mathbf{x}^{\ast}.$
+On the other hand, since $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$, the basic feasible solution $\mathbf{x}^{\ast}$ (with nonbasic variables equal to zero) is feasible for the restricted subproblem as well.
+By optimality of $\mathbf{x}^{k+1}$ for the restricted subproblem, we have
+$c^{\top}\mathbf{x}^{k+1}\ \leq\ c^{\top}\mathbf{x}^{\ast}.$
+Combining the two inequalities yields $c^{\top}\mathbf{x}^{k+1}=c^{\top}\mathbf{x}^{\ast}$, so $\mathbf{x}^{k+1}$ attains the global optimal value of the full OT problem.
+Therefore $\mathbf{x}^{k+1}$ is globally optimal.
+(ii) Assume $\mathbf{x}^{\ast}$ is the unique globally optimal solution.
+Let $j\notin\mathcal{H}_{k}$ with $x^{k}_{j}>0$.
+By the definition of the restricted subproblem, the $j$-th component cannot change:
+$x^{k+1}_{j}\;=\;x^{k}_{j}+d^{k}_{j}\;=\;x^{k}_{j}\;>\;0.$
+Now suppose (for contradiction) that $\mathbf{x}^{k+1}$ is globally optimal.
+By uniqueness of the global optimizer, this would imply $\mathbf{x}^{k+1}=\mathbf{x}^{\ast}$, and in particular $x^{\ast}_{j}=x^{k+1}_{j}>0$.
+However, under the assumption $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$, we have $j\notin\mathcal{B}^{\ast}$, and since $\mathcal{B}^{\ast}$ is a basis of a basic feasible optimal solution, all variables outside $\mathcal{B}^{\ast}$ are nonbasic and therefore equal to zero.
+Thus $x^{\ast}_{j}=0$, contradicting $x^{\ast}_{j}>0$.
+Consequently, $\mathbf{x}^{k+1}\neq\mathbf{x}^{\ast}$, and $\mathbf{x}^{k+1}$ cannot be globally optimal.
 ∎
-The above analysis motivates a practical strategy, which we call basis variable succession: at each iteration, we selecte the block $\mathcal{H}_{k}$ includes the basis variables $\mathcal{B}_{k}$ obtained from the previous subproblem.
-This design contrasts with existing BCD approaches [xie2023randomized] that select blocks randomly.
-we can solve the subproblem in iteration $k$ with complexity $O(L_{k}|\mathcal{H}_{k}|)$, which is significantly lower than the $O(LN)$ complexity of solving the full problem directly, where $L_{k}$ and $L$ denote the number of NS iterations for the subproblem and the full problem, respectively.
-This strategy offers two practical advantages:
-  1. 1.
-    The subproblem always includes all constraints and at least $M-1$ variables, ensuring feasibility.
-  2. 2.
-    Reusing the previous basis variables as the initialization eliminates the need for recomputing a starting basis in the NS method.
+The above result motivates a deterministic working set construction rule, which we call basis-variable succession.
+Specifically, at each iteration $k$, we choose the block $\mathcal{H}_{k}$ so that it contains the current basis, i.e., $\mathcal{B}_{k}\subseteq\mathcal{H}_{k}$.
+This choice contrasts with existing BCD-type approaches [18], where the working set is generated without explicitly preserving the current basis.
+From a computational perspective, let $L_{k}$ denote the number of NS pivots required to solve the restricted subproblem over $\mathcal{H}_{k}$, and let $L$ denote the pivot count when applying the NS method to the full problem.
+Then the dominant cost of the restricted solve scales as $O(L_{k}\,|\mathcal{H}_{k}|)$, which can be substantially smaller than the $O(LN)$ cost incurred by scanning all $N$ variables in each pivot of the full problem.
+This strategy has two immediate practical benefits.
+First, since the restricted subproblem retains all equality constraints and contains at least M1M-1 variables, it admits a feasible spanning tree basis and hence remains well-posed throughout the iterations.
+Second, reusing $\mathcal{B}_{k}$ as a warm start provides a readily available initial basis for the next subproblem, avoiding the overhead of repeatedly constructing an initial feasible basis from scratch.
 4.2 A Prototype of Block Coordinate Descent Network Simplex Methods and its Finite Termination
-Following the strategy of §4.1, choose a block
+Following the strategy of section 4.1, choose a block
 $\mathcal{H}_{k}\supseteq\mathcal{B}_{k},\qquad\lvert\mathcal{H}_{k}\rvert\geq M-1,$
@@
 If $\mathcal{N}_{k}=\varnothing$, then $\bar{c}^{\,k}\geq 0$ and, by complementary slackness, $\mathbf{x}^{k}$ is already globally optimal.
-Under condition (6) each subproblem produces a strictly improving search direction, ensuring monotone decrease of the objective value.
+Let $q(\mathbf{x}^{k};\mathcal{H}_{k})$ denote the optimal value of the restricted subproblem $\mathcal{D}(\mathbf{x}^{k};\mathcal{H}_{k})$.
 The next theorem establishes finite termination and global optimality of the resulting sequence $\{\mathbf{x}^{k}\}$.
 Theorem 3.
-The proposed prototype of BCDNS method satisfies the following properties:
-  1. 1.
-    Monotonicity and Convergence: The objective function decreases monotonically at each iteration, and the method converges to the global optimal solution.
-  2. 2.
-    Finite Termination: The method terminates within a finite number of iterations.
-Proof of Part 1.
-Monotonicity: For any iteration $k$, based on the selection rule for $\mathcal{I}_{k}$, there exists at least one $l\in\mathcal{I}_{k}$ such that the reduced cost of $x_{l}$ is negative.
-By the properties of the reduced cost, adding $x_{l}$ to the basis variables decreases the total cost, i.e.,
-$q(\mathbf{x}^{k};\mathcal{B}_{k}\cup\{l\})<q(\mathbf{x}^{k};\mathcal{B}_{k}).$
-where $q(\mathbf{x}^{k};\mathcal{I}_{k})$ is defined by (4) as the optimal value of the subproblem associated with the index set $\mathcal{I}_{k}$. Since $l\in\mathcal{I}_{k}$, it follows that:
-$q(\mathbf{x}^{k};\mathcal{H}_{k})=q(\mathbf{x}^{k};\mathcal{B}_{k}\cup\mathcal{I}_{k})\leq q(\mathbf{x}^{k};\mathcal{B}_{k}\cup\{l\})<q(\mathbf{x}^{k};\mathcal{B}_{k}).$
-Furthermore, since $\mathcal{J}_{k}$ is the optimal basis variables of the previous subproblem, we have:
-$q(\mathbf{x}^{k};\mathcal{B}_{k})=q(\mathbf{x}^{k-1};\mathcal{H}_{k-1}).$
-Combining the above, it follows that:
-$q(\mathbf{x}^{k};\mathcal{H}_{k})<q(\mathbf{x}^{k-1};\mathcal{H}_{k-1}).$
-Thus, the cost function decreases monotonically at each iteration.
-Global Optimality: After solving each subproblem $\mathcal{D}(\mathbf{x}^{k};\mathcal{H}_{k})$, we compute the reduced costs of all variables.
-If all reduced costs satisfy $r_{ij}\geq 0$, the algorithm terminates; otherwise, the next iteration begins.
-By the properties of the reduced cost, this condition guarantees that the final solution is globally optimal.
-∎
-Proof of Part 2.
-The total number of possible combinations of the basis variables among $N$ variables is at most $\binom{N}{M-1}$.
-Based on the selection rule, in each iteration, the set $\mathcal{I}_{k}$ always contains at least one variable with a negative reduced cost.
-Denote the variable with the smallest reduced cost as $x_{l}$.
-When solving the subproblem using the NS method, $x_{l}$ will be added to the basis variables, while one of the current basis variables $\mathcal{B}_{k}$will leave the basis.
-This implies that each iteration examines at least one new combination of the basis variables.
-Consequently, after at most $\binom{N}{M-1}$ iterations, all possible combinations of the basis variables will have been examined.
-Therefore, the algorithm is guaranteed to terminate within a finite number of iterations.
-∎
+Consider the prototype BCDNS method with the block selection rule (7).
+The sequence $\{q(\mathbf{x}^{k};\mathcal{H}_{k})\}_{k\geq 0}$ is non-increasing monotone.
+Moreover, if the NS subproblem solver is equipped with Bland’s anti-cycling pivot rule [4], then the method after finitely many degenerate pivots, either a strictly improving pivot occurs or the method terminates.
+In particular, the algorithm terminates in finitely many iterations and returns a globally optimal OT solution.
+Proof.
+Fix an outer iteration $k$ and consider the working set $\mathcal{H}_{k}=\mathcal{B}_{k}\cup\mathcal{I}_{k}$.
+We solve the restricted subproblem $\mathcal{D}(\mathbf{x}^{k};\mathcal{H}_{k})$ by the NS method started on the current basis $\mathcal{B}_{k}$.
+Monotonicity (non-increase).
+A basic property of the NS method is that each pivot maintains feasibility and does not increase the objective value; hence the optimal value of the restricted subproblem satisfies
+$q(\mathbf{x}^{k};\mathcal{H}_{k})\;\leq\;q(\mathbf{x}^{k};\mathcal{B}_{k}).$ (8)
+Since $\mathcal{B}_{k}$ is the optimal basis obtained from the previous restricted subproblem, the start objective equals the previous subproblem optimum:
+$q(\mathbf{x}^{k};\mathcal{B}_{k})\;=\;q(\mathbf{x}^{k-1};\mathcal{H}_{k-1}).$ (9)
+Combining (8)–(9) yields
+$q(\mathbf{x}^{k};\mathcal{H}_{k})\;\leq\;q(\mathbf{x}^{k-1};\mathcal{H}_{k-1}),$
+so $\{q(\mathbf{x}^{k};\mathcal{H}_{k})\}$ is non-increasing monotone.
+Finite termination and eventual strict decrease under Bland’s rule.
+Degenerate pivots may yield zero objective improvement, so strict decrease cannot be guaranteed at every pivot.
+However, when Bland’s anti-cycling pivot rule is used in the NS solver [4], cycling is prevented and hence a basis cannot repeat.
+Since the number of feasible bases is finite, each restricted NS solve terminates in finitely many pivots; moreover, it cannot perform infinitely many zero-improvement pivots, i.e., after finitely many degenerate pivots it either makes a strictly improving pivot or reaches restricted optimality.
+At the outer iteration, if the negative reduced cost set $\mathcal{N}_{k}$ is empty, then all reduced costs are nonnegative and $\mathbf{x}^{k}$ is globally optimal, so the method terminates.
+Otherwise, $\mathcal{N}_{k}\neq\varnothing$ and the block selection rule (7) guarantees that the restricted subproblem contains at least one negative reduced cost index; therefore the NS solver of the restricted subproblem at iteration $k$ is well defined and terminates after finitely many pivots.
+Consequently, the overall method terminates in finitely many outer iterations.
+4.3 Random Selection Block Coordinate Descent Network Simplex Method
+A natural way to instantiate the prototype in section 4.2 is to construct the working set by random sampling.
+we construct at iteration $k$ a working set $\mathcal{H}_{k}\supseteq\mathcal{B}_{k}$ by augmenting the current basis $\mathcal{B}_{k}$ with a randomly sampled subset $\mathcal{I}_{k}$.
+The key requirement inherited from (7) is that the block must contain at least one index with negative reduced cost whenever the current iterate is not yet globally optimal.
+Fix a sampling ratio $s\in(0,1)$ and set $b:=\lvert\mathcal{B}_{k}\rvert=M-1$.
+At iteration $k$, we draw $\lvert\mathcal{I}_{k}\rvert=\lceil sN\rceil$ indices uniformly at random from the complement of the current basis, $\{1,\dots,N\}\setminus\mathcal{B}_{k}$.
+We compute the reduced costs of the sampled indices and enforce the descent condition
+$\mathcal{I}_{k}\cap\mathcal{N}_{k}\neq\varnothing,\qquad\mathcal{N}_{k}:=\{\ell:\bar{c}^{k}_{\ell}<0\}.$
+Then we form
+$\mathcal{H}_{k}:=\mathcal{B}_{k}\cup\mathcal{I}_{k}\quad\text{and solve}\quad\mathcal{D}(\mathbf{x}^{k};\mathcal{H}_{k})\ \text{by the NS method.}$
+The algorithm is presented below.
+Algorithm 1 Random Selection Block Coordinate Descent Network Simplex Method (RS-BCDNS)
+Initialization: Find an initial feasible solution $\mathbf{x}^{0}$ and an initial basis $\mathcal{B}_{0}$ Choose a sampling ratio $s\in(0,1)$.
+for $k=0,1,2,\dots$ do
+  Compute reduced costs for all indices. If $\bar{c}^{k}_{\ell}\geq 0$ for all $\ell=1,\dots,N$ terminate
+  repeat
+   Sample $\mathcal{I}_{k}\subseteq\{1,\dots,N\}\setminus\mathcal{B}_{k}$ uniformly at random with $\lvert\mathcal{I}_{k}\rvert=\lceil sN\rceil$.
+   Compute $\bar{c}^{k}_{\ell}$ for all $\ell\in\mathcal{I}_{k}$.
+  until $\min_{\ell\in\mathcal{I}_{k}}\bar{c}^{k}_{\ell}<0$
+  Set $\mathcal{H}_{k}\leftarrow\mathcal{B}_{k}\cup\mathcal{I}_{k}$.
+  Solve $\mathcal{D}(\mathbf{x}^{k};\mathcal{H}_{k})$ by the NS method from $\mathcal{B}_{k}$, obtaining $d^{k}$ and the updated basis $\mathcal{B}_{k+1}$.
+  Update $\mathbf{x}^{k+1}\leftarrow\mathbf{x}^{k}+d^{k}$.
+end for
 4.4 Grouped Selection Block Coordinate Descent Network Simplex method
-Since an optimal solution to the OT problem is typically sparse, most variables take value zero and do not belong to the optimal basis.
-Consequently, selecting $\mathcal{I}_{k}$ uniformly at random often results in the inclusion of irrelevant variables, slowing convergence.
-To address this issue, we propose a grouped block coordinate selection strategy that partitions the variable index set $\{1,\dots,N\}$ into two disjoint subsets:
-  • $\mathcal{R}$: indices of variables deemed likely to enter the optimal basis $\mathcal{B}^{*}$,
-  • $\mathcal{S}$: indices of variables with low likelihood of being part of the optimal basis $\mathcal{B}^{*}$.
+In section 4.3, we introduced a random selection variant as a natural baseline.
+Although this random augmentation is simple and preserves the basis-succession mechanism, it can be highly inefficient for large-scale OT instances.
+Indeed, optimal transport plans are typically sparse, and only a small fraction of variables eventually belong to an optimal basis.
+Consequently, sampling $\mathcal{I}_{k}$ uniformly at random tends to include many variables that are unlikely to become basis, so their reduced costs are evaluated repeatedly and they may participate in subproblems without contributing to meaningful progress.
+To this end, we propose a grouped block-coordinate selection strategy that partitions the index set $\{1,\dots,N\}$ into two disjoint subsets:
+  • $\mathcal{R}$: indices that are deemed likely to enter an optimal basis $\mathcal{B}^{\ast}$,
+  • $\mathcal{S}$: indices with low likelihood of being part of $\mathcal{B}^{\ast}$.
 Let $0<s<t\leq 1$ be fixed sampling parameters.
@@
 Choose parameters $s,t\in(0,1)$ such that $s<t$.
-Initialize $\mathcal{S}=\emptyset$ and =$\mathcal{R}=\{1,2,\ldots,N\}$.
+Initialize $\mathcal{S}=\emptyset$ and $\mathcal{R}=\{1,2,\ldots,N\}$.
 Shuffle $\mathcal{R}$ randomly.
@@
 If all selected variables have non-negative reduced costs, move them to $\mathcal{S}$ and repeat Step 1.
-     Step 2: Move the non-selected variables with positive reduced costs to $\mathcal{S}$, and return variables with non-positive reduced costs to $\mathcal{R}$.
+  Step 2: Move the non-selected variables to $\mathcal{S}$.
   Step 3: Set $\mathcal{H}_{k}=\mathcal{I}_{k}\cup\mathcal{B}_{k}$.
@@
   Step 5: Update $x^{k+1}:=x^{k}+d^{k}$.
-     Step 6: Optimality check: If $\mathcal{R}=\emptyset$, compute reduced costs for all variables. If $r_{ij}\geq 0$ for all $i,j$, terminate; otherwise, return to Step 1.
+  Step 6: Optimality check: If $\mathcal{R}=\emptyset$, compute reduced costs for all variables.
+  if $r_{ij}\geq 0$ for all $i,j$ then
+   terminate.
+  else
+   Update $\mathcal{R}\leftarrow\{(i,j):r_{ij}<0\}$.
+   return to Step 1.
+  end if
 end for
-Let $\tilde{k}$ denote the total number of outer iterations, and let $l_{k}$ represent the number of NS iterations required to solve the subproblem at iteration $k$.
+Let $\tilde{k}$ denote the total number of iterations, and let $l_{k}$ represent the number of NS iterations required to solve the subproblem at iteration $k$.
 The time complexity of the GS-BCDNS method is given by:
@@
 Compared to the Sinkhorn algorithm, which requires $O(N)$ space, the GS-BCDNS method is more memory-efficient.
-However, relative to existing BCD methods, the memory usage depends on the size of the selected block $\mathcal{H}_{k}$ [xie2023randomized].
 5.  Numerical Experiments
-In this section, we perform numerical experiments on various instances of OT problems1.11footnotetext: 1 Numerical experiments were conducted on a computer with an Intel Core i7-12700H CPU (2.70 GHz) and 16 GB of RAM.
-The implementation was performed using Python 3.8, and the code has been made publicly available on GitHub at https://github.com/842422492/BCDNS.git.
-We first solve randomly generated OT problems using the GS-BCDNS method to evaluate the impact of parameters $s$ and $t$ on algorithmic efficiency.
-Subsequently, we compute the Wasserstein distance between uniform and normal distributions using different algorithms to compare their performance.
-5.1 Experiment Settings
-For simplicity, we assume that the numbers of supply and demand nodes are equal, i.e., $n=m$.
-Test Problem 1: Ten randomly generated discrete OT problems using randomly sampled numbers. $n=1000$.
-Test Problem 2: The Wasserstein distance between a one-dimensional uniform distribution on $[-1,1]$ and a standard normal distribution on $\mathbb{R}$. $n=1000$.
-Test Problem 3: The Wasserstein distance between a two-dimensional uniform distribution on $[-1,1]^{2}$ and a standard normal distribution on $\mathbb{R}^{2}$. $n=1000$.
-Test Problem 4: The Wasserstein distance between a three-dimensional uniform distribution on $[-1,1]^{3}$ and a standard normal distribution on $\mathbb{R}^{3}$. $n=1000$.
-Test Problem 5: The Wasserstein distance between a one-dimensional uniform distribution on $[-1,1]$ and a standard normal distribution for different problem sizes.
-$n$ starts from 50 and increases by 50 up to 300.
-The discrete OT problems in Test Problem 1 were generated as follows: two sets of $n=1000$ random numbers were sampled from a uniform distribution on $[1,1000]$ to represent the supply and demand distributions $P$ and $Q$, respectively.
-Then adjuste the two sets to ensure that their sums were equal.
-Additionally, a cost matrix of size $1000\times 1000$ was generated from the same uniform distribution.
-The Wasserstein distances in Test Problem 2, 3, and 4 were estimated as follows: $n$ samples were randomly drawn from the uniform and normal distributions, denoted as $y$ and $\tilde{y}$, respectively.
-For each pair, i.i.d.
-samples $\{\tilde{y}_{i}\}$ and $\{y_{j}\}$, $1\leq i,j\leq 1000$, were used to approximate the two continuous probability measures.
-The cost matrix was defined as: $c_{i,j}=\text{dist}(\tilde{y}_{i},y_{j})^{2},$ and the probability mass vectors were set as $\mathbf{p}=\mathbf{q}=(1/n,\dots,1/n)^{T}\in\mathbb{R}^{n}$.
-Figure 1: The point distributions in Test Problem 2, 3 and 4
-Algorithm Settings:
-  • Sinkhorn (Stabilized) Algorithm: As used in [xie2023randomized], we follow the method proposed in [cuturi2013sinkhorn] for computing the Wasserstein distance.
-Let $\gamma$ be the entropy regularization coefficient, which is set to $\gamma=1/(4\log n)$ as suggested in [dvurechensky2018computational].
-We consider the parameter settings $\varepsilon=10^{-4},10^{-3},0.01,0.1$.
-Sinkhorn iterations are projected onto the feasible region using a rounding procedure, specifically Algorithm 2 in [altschuler2017near].
-To mitigate numerical instability, all updates are performed in log space using the LogSumExp function.
-The algorithm terminates when either the change in the objective function value falls below $\delta=10^{-6}$ or the runtime exceeds 2000 seconds.
-  • GS-BCDNS: When solving Test Problem 1, the parameter $s$ was varied from $1/n$ to 0.9, while the parameter $t$ was varied from $2/n$ to 1.
-For Test Problems 2, 3, 4 and 5, the values of $s$ and $t$ were fixed at $5/n$ and $10/n$, respectively.
-For each Test Problem, set $\{e_{k}\}$ to positive infinity and calculate the initial basic feasible solution using the Northwest Corner Rule.
+In this section, we report numerical experiments designed to (i) study the effect of the block-size parameters $(s,t)$ in the proposed BCDNS framework, (ii) compare the proposed methods with representative baselines on the Wasserstein distance computations, and (iii) evaluate the performance and scalability of the proposed methods on large-scale OT instances.11footnotetext: Numerical experiments were conducted on a computer with an Intel Core i7-12700H CPU (2.70 GHz) and 16 GB of RAM.
+The implementation was performed in Python 3.8.
+The code and the scripts for reproducing all figures are publicly available at https://github.com/842422492/BCDNS.git.
+5.1 Experimental Setup
+This section describes the common procedure used to generate OT test instances and the implementation details shared by all methods in Section 5.2–Section 5.4.
+Unless otherwise stated, we consider balanced OT with equal numbers of supply and demand nodes, i.e., $n=m$.
+Test instance generation and reproducibility.
+All Wasserstein test instances are constructed by Monte Carlo discretization of continuous probability measures.
+For a prescribed sample size $n$, we draw i.i.d.
+samples $\{u_{i}\}_{i=1}^{n}$ from a source distribution $\mu$ and $\{v_{j}\}_{j=1}^{n}$ from a target distribution $\nu$.
+We define the quadratic transport cost
+$c_{ij}\;=\;\|u_{i}-v_{j}\|^{2},\qquad 1\leq i,j\leq n,$ (10)
+and set the discrete marginals to be uniform,
+$p=q=\Bigl(\tfrac{1}{n},\dots,\tfrac{1}{n}\Bigr)^{\top}\in\mathbb{R}^{n}.$ (11)
+The resulting linear program is then solved by the competing methods.
+To ensure reproducibility, we save and version-control all sampled point sets $\{u_{i}\}$ and $\{v_{j}\}$ for every reported instance; hence the corresponding cost matrix $C$ in (10) can be reconstructed exactly.
+Algorithmic settings.
+  • Sinkhorn (stabilized).
+Following [6] and the stabilized implementation used in [18], we solve the entropically regularized OT problem with regularization coefficient $\gamma=\varepsilon/(4\log n)$ as suggested in [7].
+We test $\varepsilon\in\{10^{-4},10^{-3},10^{-2},10^{-1}\}$.
+Each Sinkhorn iterate is rounded to feasibility using the projection/rounding procedure (Algorithm 2 in [2]).
+To improve numerical stability, all updates are carried out in log-domain via the LogSumExp operator.
+The algorithm terminates when the absolute change in objective value falls below $\delta=10^{-6}$ or when the runtime exceeds $2000$ seconds.
+  • GS-BCDNS.
+    Our NS method is implemented by following the design of the network simplex solver in the Python NetworkX package.
+In the experiments, we multiply the cost matrix by a positive scaling factor and round to integers to obtain an equivalent integer-cost instance for the NS pivots; this scaling does not change the set of optimal transport plans, since it preserves the ordering of feasible solutions by objective value.
+In Section 5.2, we first perform a search on $(s,t)$ to identify an efficient regime for subproblem sizes.
+Based on the empirical findings in Section 5.2, we fix $(s,t)$ for all subsequent comparisons.
+Unless otherwise stated, we use
+    $s=\tfrac{2}{n},\qquad t=\tfrac{20}{n}.$
+  • RS-BCDNS. The method is initialized by the Northwest Corner method and set the block size to 256.
 5.2 Effect of Parameters $s$ and $t$
-In the BCDNS method, the parameters $s$ and $t$ play a crucial role.
-Here, we vary the parameters $s$ and $t$ to solve Test Problem 1.
-The number of reduced cost computations serves as a key indicator in the BCDNS method.
-For each case, we record the corresponding number of pivots and reduced cost computations.
-And then we calculate the averages of these metrics to evaluate the algorithm’s performance.
-In this context, the number of pivots refers to the number of iterations performed by the NS method.
-Figure 2: Effects of $s$ and $t$ on the number of pivots and reduced cost computations
-Figure 1 illustrates the effects of $s$ and $t$ on two key performance metrics: the number of pivots and the total number of reduced cost computations.
-We can observe that the number of pivots decreases as $s$ increases, as larger subproblem sizes reduce the total iterations required by the NS method.
-However, as $s$ increases, the size of each subproblem also grows, which, while reducing the number of pivots, increases the computational cost of each pivot.
-Similarly, increasing $t$ enhances the quality of the variables selected for the subproblem, further reducing the number of pivots.
-For small values of $s$, increasing $s$ reduces the total number of reduced cost computations by enabling more efficient solutions to larger subproblems.
-However, beyond a certain threshold, further increases in $s$ lead to an overall rise in computational effort due to the greater complexity of individual subproblems.
-The effect of $t$ is less pronounced, except when $s$ is very small, in which case increasing $t$ significantly increases the computational burden, as the algorithm approaches the behavior of the standard NS method.
-Extensive numerical experiments consistently reveal that reduced cost computations are minimized when $t$ is approximately $2s$.
+In GS-BCDNS, the sampling parameters $s,t\in(0,1)$ determine the working-set size $\lvert\mathcal{H}_{k}\rvert\approx(M-1)+sN$ and the number of candidate variables whose reduced costs are inspected before forming $\mathcal{I}_{k}$.
+This subsection investigates how $(s,t)$ affects computational cost in practice.
+Test Problem 1. (Uniform–Normal, varying $s$ and $t$)
+We compute the Wasserstein distance between the uniform distribution on [1,1][-1,1] and the standard normal distribution $\mathcal{N}(0,1)$ by drawing $x_{i}\sim\mathrm{Unif}[-1,1]$ and $y_{j}\sim\mathcal{N}(0,1)$ using $i.i.d.$ samples of equal size $n\in\{250,500,1000\}$.
+For each pair $(s,t)$, we run GS-BCDNS until the optimality test is satisfied and record: (i) the total number of reduced-cost evaluations (denoted “reduced-cost evals”), (ii) the total number of pivots performed by the network simplex solver across all subproblems (denoted “pivot count”), and (iii) wall-clock runtime in seconds.
+The reduced-cost evaluations measure the dominant arithmetic workload, because each evaluation is an $O(1)$ operation in the OT network structure, and they are performed repeatedly during candidate screening and optimality checks.
+Comment on Test Problem 1.
+Figures 13 visualize the three metrics as heatmaps over $(s,t)$. The white regions correspond to parameter pairs that were not evaluated (in particular, those that violated $s<t$).
+Across all $n$, the runtime heatmap closely matches the reduced-cost evaluation heatmap, indicating a strong positive correlation between these two quantities: configurations that trigger more reduced-cost computations consistently take longer in wall-clock time.
+Moreover, for all three problem sizes, both reduced-cost evaluations and runtime attain their smallest values when $s$ is chosen on the order of $2/n$ – -$3/n$ and $t$ is on the order of $20/n$ – -$30/n$, i.e.,
+$s\in\Bigl[\tfrac{2}{n},\,\tfrac{3}{n}\Bigr],\qquad t\in\Bigl[\tfrac{20}{n},\,\tfrac{30}{n}\Bigr].$
+suggesting that a small working-set fraction together with a moderate candidate screening fraction offers the best trade-off between (a) selecting sufficiently good entering candidates and (b) avoiding excessive reduced-cost screening overhead.
+In contrast, the pivot count depends much more strongly on $s$ than on $t$.
+As $s$ increases, the subproblem size grows and the network simplex solver needs fewer pivots to achieve the optimality of the subproblem, leading to a clear decrease in pivot count.
+However, a larger $s$ also increases the per-iteration burden through a larger $\lvert\mathcal{H}_{k}\rvert$ and more bookkeeping within each subproblem; consequently, beyond a certain threshold, the reduced-cost workload increases again even though the pivot count continues to fall.
+This shows that the pivot count alone is not a reliable proxy for the total runtime.
+In general, qualitative patterns are consistent for $n=250,500,1000$, while absolute magnitudes increase with problem size as expected.
+Figure 1: Effect of $(s,t)$ on reduced-cost evaluations, pivot count, and runtime for the instance with $n=250$.
+Figure 2: Effect of $(s,t)$ on reduced-cost evaluations, pivot count, and runtime for the instance with $n=500$.
+Figure 3: Effect of $(s,t)$ on reduced-cost evaluations, pivot count, and runtime for the instance with $n=1000$.
 5.3 Comparison with the Existing Methods
-In this chapter, we primarily evaluate the performance differences among the GS-BCDNS method, the Sinkhorn algorithm, and the NS method.
-Figure 3: Comparison of algorithms to computer Wasserstein distance
-As described in Section 4.1, we apply the Sinkhorn algorithm with different values of $\epsilon$ and the GS-BCDNS method to solve Test Problems 2, 3 and 4.
-The stopping criterion for the Sinkhorn algorithm is set such that the algorithm terminates when the change in the objective function value falls below $\delta=10^{-6}$ or when the runtime exceeds 2000 seconds.
-The GS-BCDNS method is tested with fixed parameters $s=5/n$ and $t=10/n$.
-In the evaluation, the x-axis represents the runtime, while the y-axis represents the gap between the current objective function value and the optimal objective function value, where the optimal solution is obtained using the interior-point method implemented in Python’s SciPy library.
-The results in Figure 3 indicate that for larger values of $\epsilon$ (i.e, $\epsilon=0.1$ and $\epsilon=0.01$), the Sinkhorn algorithm exhibits a faster convergence speed than the GS-BCDNS method; however, the approximation error remains relatively large.
-When $\epsilon=0.001$, the convergence speed of the Sinkhorn algorithm is comparable to that of the GS-BCDNS method, but its final convergence is slightly slower.
-For extremely small values of $\epsilon$, the convergence speed of the Sinkhorn algorithm is significantly slower than that of the GS-BCDNS method.
-Figure 4: Comparison of algorithms to solve Test Problem 5
-Figure 4 presents the results of solving Test Problem 5 using both the NS method and the GS-BCDNS method.
-The y-axis of Figure 4 represents the ratio of the NS method’s runtime to that of the GS-BCDNS method.
-The results demonstrate that the GS-BCDNS method significantly outperforms the NS method across all tested datasets.
+In this subsection, we compare the proposed GS-BCDNS with the existing method.
+First, we compare the proposed GS-BCDNS with the classical NS method and the baseline RS-BCDNS method on Test Problem 2 and 3.
+We also report (i) the wall-clock runtime and (ii) the total number of reduced-cost evaluations.
+Consistent with this, the experiments show a strong correlation between the reduced-cost evaluations and the runtime.
+Second, We compare GS-BCDNS with the stabilized Sinkhorn algorithm on Test Problem 4.
+For Sinkhorn, we test multiple regularization levels $\varepsilon\in\{10^{-1},10^{-2},10^{-3},10^{-4}\}$ under the same stabilization and stopping rule as in the previous experiments.
+For evaluation, we report the optimality gap $\mathrm{Gap}(t)=f(x(t))-f^{*}$ as a function of wall-clock time.
+Test Problem 2 (Uniform–Normal, varying $n$).
+We compute the Wasserstein distance between the uniform distribution on [1,1][-1,1] and the standard normal distribution $\mathcal{N}(0,1)$ by drawing $x_{i}\sim\mathrm{Unif}[-1,1]$ and $y_{j}\sim\mathcal{N}(0,1)$.
+We vary the problem size as $n\in\{50,100,150,200,250,300,350,400\}$.
+Test Problem 3 (Normal–Mixture, varying $n$).
+We compute the Wasserstein distance between $\mathcal{N}(0,1)$ and a two-component Gaussian mixture
+$\frac{1}{2}\,\mathcal{N}(-2,1)\;+\;\frac{1}{2}\,\mathcal{N}(2,1),$
+by sampling $x_{i}\sim\mathcal{N}(0,1)$ and $y_{j}\sim\frac{1}{2}\mathcal{N}(-2,1)+\frac{1}{2}\mathcal{N}(2,1)$. The same range of $n$ is tested.
+Comment on Test Problem 2.
+Figure 4 shows that GS-BCDNS consistently achieves the smallest runtime and the fewest reduced-cost evaluations over all tested $n$.
+Moreover, the advantage becomes more pronounced as $n$ increases: GS-BCDNS reduces the number of reduced-cost evaluations by roughly an order of magnitude compared with NS.
+In contrast, RS-BCDNS provides only modest improvements over NS.
+The corresponding speedup curves in Figure 5 indicate that GS-BCDNS attains large and increasing speedups with respect to NS as $n$ grows, whereas RS-BCDNS stays close to the baseline.
+Comment on Test Problem 3.
+The Normal–Mixture instances are empirically more challenging for random block construction.
+As shown in Figure 6, GS-BCDNS remains the best-performing method across all tested sizes, again substantially reducing reduced-cost evaluations and runtime.
+However, RS-BCDNS can become slower than NS for medium-to-large $n$.
+This phenomenon is consistent with the mechanism of RS-BCDNS: random augmentation tends to include many variables that are unlikely to belong to the final optimal basis, leading to repeated subproblem solves with limited progress while still incurring significant reduced-cost computations.
+The speedup plot in Figure 7 confirms that RS-BCDNS may fall below the NS baseline, whereas GS-BCDNS maintains a clear advantage.
+Figure 4: Test Problem 2: runtime (left) and total reduced-cost evaluations (right) for NS, RS-BCDNS, and GS-BCDNS.
+Figure 5: Test Problem 2: speedup relative to NS (NS time / method time).
+Figure 6: Test Problem 3: runtime (left) and total reduced-cost evaluations (right) for NS, RS-BCDNS, and GS-BCDNS.
+Figure 7: Test Problem 3: speedup relative to NS (NS time / method time).
+Test Problem 4 (1D/2D Uniform–Normal).
+Wasserstein distance estimation between (i) a one-dimensional uniform distribution on [1,1][-1,1] and a standard normal distribution, and (ii) their two-dimensional counterparts on $1,1]^{2}$ and $\mathcal{N}(0,I_{2})$.
+Comment on Test Problem 4.
+Figures 89 show a consistent accuracy–speed trade-off.
+With relatively large regularization, Sinkhorn quickly reduces the objective at the beginning, but the gap stagnates at a non-negligible level due to the inherent regularization bias.
+As $\varepsilon$ decreases, Sinkhorn becomes progressively closer to the unregularized OT solution, yet its convergence speed deteriorates substantially.
+In contrast, GS-BCDNS exhibits a slower initial decrease than low-precision Sinkhorn, but it continues to make steady progress and ultimately attains the exact optimal solution of the original OT problem.
+Figure 8: Test Problem 4 (1D): optimality gap versus time for GS-BCDNS and stabilized Sinkhorn under different regularization levels $\varepsilon$.
+Figure 9: Test Problem 4 (2D): optimality gap versus time for GS-BCDNS and stabilized Sinkhorn under different regularization levels $\varepsilon$.
+5.4 Large-scale Experiment
+This subsection evaluates the proposed method on a larger OT instance with $n=4000$, where solving the full problem directly becomes significantly more demanding.
+We consider a one-dimensional Wasserstein distance problem between a uniform source distribution and a heavy-tailed target distribution.
+Test Problem 5 (Uniform–Beta).
+The source samples $\{u_{i}\}_{i=1}^{n}$ are drawn from $\mathrm{Unif}([-1,1])$.
+The target samples $\{v_{j}\}_{j=1}^{n}$ are drawn from a $\mathrm{Beta}(0.5,0.5)$ distribution and then linearly mapped to [1,1][-1,1].
+To interpret the evolution of the computed transport plan, we use the barycentric projection
+$T_{k}(u_{i})\;:=\;\frac{1}{p_{i}}\sum_{j=1}^{n}v_{j}\,x^{k}_{ij},\qquad i=1,\dots,n,$
+where $x^{k}$ denotes the transport plan produced after $k$ iterations. Intuitively, $T_{k}(u_{i})$ is the average target location to which the mass in $u_{i}$ is transported under $x^{k}$.
+Comment on Test Problem 5.
+Figure 10 summarizes the large-scale experiment.
+The top panel plots the empirical marginals: the source is approximately uniform, while the target (Beta(0.5,0.5)(0.5,0.5) mapped to [1,1][-1,1]) concentrates more mass near the two endpoints, producing a characteristic U-shaped density.
+This structural mismatch implies that the optimal coupling must transport a nontrivial portion of the mass from the middle region of the uniform source toward the two extremes of the target support.
+The bottom panel shows the barycentric projection curves $T_{k}(x)$ at several epochs ($k=10,20,50,100$) together with the final optimal map (“Optimal”).
+At early epochs (e.g.
+$k=10$ and $k=20$), the barycentric projection exhibits substantial oscillations, reflecting that the current sparse basic solution still contains non-optimal variables and the coupling has not yet stabilized.
+As the iterations proceed, the projection becomes increasingly structured and approaches a smooth monotone map: by $k=50$ the curve already aligns closely with the optimal trend, and by $k=100$ it nearly overlaps the optimal map over most of the domain.
+Consistent with this visual evidence, the proposed method reaches the globally optimal solution at $218$ epochs in this experiment.
+In general, Figure 10 demonstrates that even at $n=4000$, the iterations produced by GS-BCDNS steadily reduce the optimality gap and the associated transport map progressively converges toward the exact optimal coupling.
+Figure 10: Large-scale OT experiment with $n=4000$.
+Top: empirical source and target marginals (Uniform[1,1][-1,1] vs.
+Beta(0.5,0.5)(0.5,0.5) mapped to [1,1][-1,1]).
+Bottom: barycentric projection $T_{k}(x)$ at selected epochs $k$, illustrating progressive convergence of the transport map toward the optimal monotone coupling.
 6.  Conclusion
-In this paper, we review the existing methods for solving the OT problems, including the NS method, the BCD method, and the Sinkhorn algorithm.
-Then we propose the BCDNS methods and explain thier theoretical foundations and the basis variable succession strategy.
-Furthermore, we provide a detailed algorithmic framework and established theoretical guarantees for finite termination and convergence to the exact optimal solution.
-Extensive numerical experiments demonstrate that the BCDNS method achieves superior performance in terms of runtime and computational cost compared to the NS method.
-While the BCDNS method converges slightly slower than the Sinkhorn algorithm with large values of $\epsilon$, it has the advantage of guaranteeing convergence to an exact solution.
-Moreover, compared to the Sinkhorn algorithm with small values of $\epsilon$, the BCDNS method exhibits even faster convergence.
-This makes the BCDNS method particularly valuable for large-scale OT problems that require exact solutions.
-Future work could explore several directions.
-First, improving the selection strategy for block coordinates may further enhance convergence speed.
-Second, extending the BCDNS method to more general network flow problems could broaden its applicability.
-Appendix A Proof of $\operatorname{rank}(A)=M-1$
-Proof.
-Let
-$A=\begin{bmatrix}I_{n}&I_{n}&\cdots&I_{n}\\ -\mathbf{1}_{n}^{\top}&0&\cdots&0\\ 0&-\mathbf{1}_{n}^{\top}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&-\mathbf{1}_{n}^{\top}\end{bmatrix}\in\mathbb{R}^{M\times N},\quad M=n+m,\;N=nm,$
-where $I_{n}$ is the $n\times n$ identity matrix and $\mathbf{1}_{n}$ is the $n$-vector of all ones.
-(i) $\operatorname{rank}(A)\leq M-1$.
-Denote the rows of $A$ by $r_{1},\dots,r_{M}$. Observe that
-[…]
-because each of the first $n$ rows contributes exactly one “$+1$” in each block and each of the last $m$ rows contributes exactly one “$-1$”, so they cancel out.
-Hence the rows are linearly dependent, and
-$\operatorname{rank}(A)\;\leq\;M-1.$
-(ii) $\operatorname{rank}(A)\geq M-1$.
-Interpret $A$ as the node–arc incidence matrix of the complete bipartite directed graph $K_{n,m}$ with $M=n+m$ nodes and $nm$ arcs.
-Choose any spanning tree $T$ of this graph, which has exactly $M-1$ edges.
-Let
-$B\;=\;A(:,E(T))\;\in\;\mathbb{R}^{M\times(M-1)}$
-be the submatrix of $A$ whose columns correspond to the edges in $T$.
-Since $T$ is a tree, its incidence matrix has rank $M-1$, and in particular there is an $(M-1)\times(M-1)$ submatrix of $B$ with determinant ±$\pm 1$.
-Therefore
-$\operatorname{rank}(A)\;\geq\;M-1.$
-Combining (i) and (ii) gives
-$\operatorname{rank}(A)\;=\;M-1.$
-[…]
+In this paper, we studied efficient algorithms to solve the OT problem.
+After reviewing representative approaches, including the NS method, BCD schemes, and the Sinkhorn algorithm, we proposed the BCDNS method.
+The proposed framework integrates NS pivots with block coordinate updates and introduces the basis variable succession strategy to preserve feasibility across subproblems.
+We established theoretical guarantees showing that the BCDNS method terminates in a finite number of iterations and converges to an exact optimal solution.
+Extensive numerical experiments demonstrate that the BCDNS method substantially improves computational efficiency over the classical NS method on large-scale instances, achieving speed-ups of up to tens of times while maintaining exact optimality.
+Compared with the Sinkhorn algorithm, the BCDNS method exhibits a different accuracy–efficiency trade-off: although it converges more slowly than Sinkhorn with large regularization parameters, it avoids regularization bias and guarantees convergence to the OT optimum.
+Moreover, when high accuracy is required, the BCDNS method can be competitive with or even faster than Sinkhorn.
+These results indicate that the BCDNS method is particularly well suited for large-scale OT problems where exact solutions are required.
+Several directions merit further investigation.
+First, more sophisticated block selection strategies may further accelerate convergence.
+Second, extending the proposed framework to more general network flow and structured linear programming problems could broaden its applicability.